Jefferson Lab HALL A Logbook for May 08 - All EXCEPT Automatic Entries

  • Main INDEX

    Saturday

  • 240591 05/31/08 23:28 A.Rakhman Target Screen Shots
  • 240585 05/31/08 20:53 M.Shabestari MCC called. No beam for a while
  • 240583 05/31/08 20:36 M.Shabestari BigBox Screens
  • 240582 05/31/08 20:24 M.Shabestari BPM & BCM stripcharts
  • 240581 05/31/08 20:16 adaq Radiation and Ozone snapshot
  • 240580 05/31/08 20:16 adaq Hall A Tools snapshot
  • 240579 05/31/08 20:15 adaq BigBox Magnet snapshot
  • 240578 05/31/08 20:15 adaq Beam Position and Current snapshot
  • 240577 05/31/08 20:15 adaq Beam Energy in A and C snapshot
  • 240576 05/31/08 20:03 jinhuang bblinux2 is down
  • 240575 05/31/08 20:03 adaq Target Screens snapshot
  • 240574 05/31/08 19:54 M.Shabestari Trippy beam
  • 240571 05/31/08 19:44 M.Shabestari Spot++ for run 1767 (LH2 6cm)
  • 240569 05/31/08 19:27 M.Shabestari BCM snapshot
  • 240567 05/31/08 18:59 Nikos 5M events/hour
  • 240566 05/31/08 18:21 jinhuang Hot Wire In FPP Fixed itself
  • 240563 05/31/08 17:32 jinhuang Hot Wire In FPP U1 Plane
  • 240562 05/31/08 17:18 A.Rakhman Target Screen Shots
  • 240561 05/31/08 17:09 jinhuang ROC 8 Reboot
  • 240557 05/31/08 16:28 M.Shabestari Hall A Tools Snapshot
  • 240556 05/31/08 16:04 Ransome Shift summary
  • 240548 05/31/08 13:30 Ransome big bite screens
  • 240547 05/31/08 13:23 Ransome spot
  • 240544 05/31/08 12:19 xiaohui Hall A Tools after Kinematics change to K8 D-2
  • 240543 05/31/08 12:15 adaq Radiation and Ozone snapshot
  • 240542 05/31/08 12:15 adaq Hall A Tools snapshot
  • 240541 05/31/08 12:15 adaq BigBox Magnet snapshot
  • 240540 05/31/08 12:15 adaq Beam Position and Current snapshot
  • 240539 05/31/08 12:15 adaq Beam Energy in A and C snapshot
  • 240538 05/31/08 12:15 xiaohui HV map K8 D-2
  • 240537 05/31/08 12:12 xiaohui Acceptance check
  • 240536 05/31/08 12:11 xiaohui BB shower blocks
  • 240535 05/31/08 12:11 Nikos # events
  • 240534 05/31/08 12:03 adaq Target Screens snapshot
  • 240530 05/31/08 11:32 ransome screen after momentum change
  • 240529 05/31/08 11:28 Nikos BigBite moved to 50 deg.
  • 240517 05/31/08 08:10 yez Target Screen(2)
  • 240516 05/31/08 08:09 yez Target Screen at the begining of Day shift(1)
  • 240513 05/31/08 08:01 M.Olson Owl Shift Summary
  • 240511 05/31/08 07:59 M.Olson Beam Current Profile (0600-0800)
  • 240508 05/31/08 06:40 M.Olson spot++ for run 1751 (LH2 6cm)
  • 240507 05/31/08 06:25 M.Olson Target Screens
  • 240506 05/31/08 06:23 M.Olson FPP Thresholds and Carbon Door Screens
  • 240505 05/31/08 06:22 M.Olson FPP HV and LV screens
  • 240504 05/31/08 06:22 M.Olson LHRS HV Screen
  • 240503 05/31/08 06:13 M.Olson BB_SH map
  • 240502 05/31/08 06:12 M.Olson BigBox Screens
  • 240501 05/31/08 06:11 M.Olson Hall A Tools
  • 240500 05/31/08 06:07 M.Olson Target to LH2 (6cm)
  • 240499 05/31/08 06:04 M.Olson Beam Current Profile (0400-0600)
  • 240496 05/31/08 05:32 M.Olson spot++ for run 1750 (Al dummy)
  • 240494 05/31/08 05:23 M.Olson Target to Al Dummy (6cm)
  • 240490 05/31/08 04:15 adaq Radiation and Ozone snapshot
  • 240489 05/31/08 04:15 adaq Hall A Tools snapshot
  • 240488 05/31/08 04:15 adaq BigBox Magnet snapshot
  • 240487 05/31/08 04:15 adaq Beam Position and Current snapshot
  • 240486 05/31/08 04:15 adaq Beam Energy in A and C snapshot
  • 240484 05/31/08 04:04 M.Olson Beam Current Profile (0200-0400)
  • 240483 05/31/08 04:03 adaq Target Screens snapshot
  • 240481 05/31/08 03:26 M.Olson Spot++ for run 1747
  • 240476 05/31/08 02:00 M. Olson Beam Current Profile (0000-0200)
  • 240475 05/31/08 01:26 M.Olson Beam back, spot++ for run 1745
  • 240473 05/31/08 01:19 M.Olson Bubbler on monitor in center room not bubbling (
  • 240472 05/31/08 01:11 M.Olson Beam down (RF) since 0030
  • Friday

  • 240469 05/30/08 23:54 Y. Zhang replay run 1744
  • 240468 05/30/08 23:52 Ransome Shift summary
  • 240463 05/30/08 22:18 xiaohui HV map K7_D+2
  • 240462 05/30/08 22:12 Y. Zhang target screen
  • 240461 05/30/08 22:12 xiaohui Acceptance check
  • 240460 05/30/08 22:09 xiaohui BB shower blocks K7_D+2
  • 240445 05/30/08 20:15 adaq Radiation and Ozone snapshot
  • 240444 05/30/08 20:15 adaq Hall A Tools snapshot
  • 240443 05/30/08 20:15 adaq BigBox Magnet snapshot
  • 240442 05/30/08 20:14 adaq Beam Position and Current snapshot
  • 240441 05/30/08 20:14 adaq Beam Energy in A and C snapshot
  • 240440 05/30/08 20:03 adaq Target Screens snapshot
  • 240426 05/30/08 15:59 Ole Hansen Day shift summary
  • 240421 05/30/08 14:24 hyao target screen
  • 240418 05/30/08 13:23 Moffit Cleaned up e04007/onlana_rootfiles on adaql10:/work1
  • 240413 05/30/08 12:14 adaq Radiation and Ozone snapshot
  • 240412 05/30/08 12:14 adaq Hall A Tools snapshot
  • 240411 05/30/08 12:14 adaq BigBox Magnet snapshot
  • 240410 05/30/08 12:14 adaq Beam Position and Current snapshot
  • 240409 05/30/08 12:14 adaq Beam Energy in A and C snapshot
  • 240408 05/30/08 12:14 xiaohui HV map K7D0
  • 240407 05/30/08 12:09 Ole Hansen Hall A Tools & BCM
  • 240406 05/30/08 12:09 xiaohui Acceptance check
  • 240405 05/30/08 12:08 xiaohui BB shower block K7D0
  • 240404 05/30/08 12:02 adaq Target Screens snapshot
  • 240403 05/30/08 12:00 Ole Hansen Lowering beam current
  • 240401 05/30/08 12:00 Ole Hansen Start production at delta = 0%
  • 240396 05/30/08 11:38 hyao target moved from Al to H2
  • 240395 05/30/08 11:32 Ole Hansen Moving to delta = 0%
  • 240392 05/30/08 11:09 hyao target moved from H2 to Al
  • 240381 05/30/08 08:02 B. Hahn Owl Shift Summarry
  • 240372 05/30/08 04:28 chiranjib SPOT ++
  • 240371 05/30/08 04:28 B. Hahn Big Bite Screens
  • 240370 05/30/08 04:27 B. Hahn Screen Shots
  • 240369 05/30/08 04:25 Chiranjib Target screens
  • 240368 05/30/08 04:14 adaq Radiation and Ozone snapshot
  • 240367 05/30/08 04:14 adaq Hall A Tools snapshot
  • 240366 05/30/08 04:14 adaq BigBox Magnet snapshot
  • 240365 05/30/08 04:14 adaq Beam Position and Current snapshot
  • 240364 05/30/08 04:14 adaq Beam Energy in A and C snapshot
  • 240363 05/30/08 04:02 adaq Target Screens snapshot
  • 240353 05/30/08 00:15 B. Hahn Big Bite HV screens
  • 240351 05/30/08 00:00 LeRose Shift Summary (swing)
  • Thursday

  • 240350 05/29/08 23:59 jinhuang Target ScreenShot
  • 240349 05/29/08 22:44 brads Run 1707 is pedestal run (ps8=1, no beam)
  • 240346 05/29/08 22:35 brads BB Cerenkov update
  • 240345 05/29/08 22:33 brads BB Cerenkov HVs turned back on.
  • 240344 05/29/08 20:14 adaq Radiation and Ozone snapshot
  • 240343 05/29/08 20:14 adaq Hall A Tools snapshot
  • 240342 05/29/08 20:14 adaq BigBox Magnet snapshot
  • 240341 05/29/08 20:14 adaq Beam Position and Current snapshot
  • 240340 05/29/08 20:14 adaq Beam Energy in A and C snapshot
  • 240339 05/29/08 20:06 LeRose Still no beam
  • 240338 05/29/08 20:02 adaq Target Screens snapshot
  • 240337 05/29/08 18:55 LeRose No Beam
  • 240332 05/29/08 17:11 xiaohui HV start from run 1704
  • 240331 05/29/08 16:58 xiaohui Start production K7-D-2 officially
  • 240330 05/29/08 16:55 xiaohui Start production K7-D-2 officially
  • 240325 05/29/08 16:25 xiaohui HV for run 1702
  • 240320 05/29/08 16:00 Holmstrom Day Shift Summary
  • 240318 05/29/08 15:43 xiaohui Acceptance check2
  • 240317 05/29/08 15:41 xiaohui Acceptance check
  • 240316 05/29/08 15:40 xiaohui BB shower blocks K7_D-2
  • 240310 05/29/08 14:57 ycwang Target moved from carbon to Loop3:6cm LH2
  • 240309 05/29/08 14:57 R. Michaels Successful plot of cavities at 50 nA
  • 240307 05/29/08 14:47 R. Michaels Cavities, part V. - Finished.
  • 240286 05/29/08 14:17 brads C4F8O fill status
  • 240285 05/29/08 14:11 R. Michaels There's hope for them cavities
  • 240284 05/29/08 13:51 R. Michaels access to install LNA
  • 240282 05/29/08 13:48 R. Michaels cavities, part IV
  • 240271 05/29/08 13:21 R. Michaels cavity test, part III.
  • 240261 05/29/08 12:58 R. Michaels cavity test, part II
  • 240253 05/29/08 12:41 brads BB Cerenkov work
  • 240252 05/29/08 12:36 R. Michaels cavity tests, part I
  • 240249 05/29/08 12:13 adaq Radiation and Ozone snapshot
  • 240248 05/29/08 12:13 adaq Hall A Tools snapshot
  • 240247 05/29/08 12:13 adaq BigBox Magnet snapshot
  • 240246 05/29/08 12:13 adaq Beam Position and Current snapshot
  • 240245 05/29/08 12:13 adaq Beam Energy in A and C snapshot
  • 240242 05/29/08 12:02 adaq Target Screens snapshot
  • 240234 05/29/08 11:20 Holmstrom Big Bite magnet on
  • 240230 05/29/08 10:22 brads Swapped out ethane/argon tank for BB MWDCs.
  • 240229 05/29/08 10:21 brads Removed bblinux1 nfs mounts from bblinux2
  • 240228 05/29/08 10:19 Moffit / All T7 Fixed, but still disconnected.
  • 240225 05/29/08 10:06 R. Michaels stripline pedestal run
  • 240221 05/29/08 09:52 ycwang Target moving from Loop3:6cm to Carbon
  • 240220 05/29/08 09:52 Holmstrom Left Arm IOC Crash
  • 240218 05/29/08 09:31 R. Michaels R-HRS work
  • 240212 05/29/08 09:25 Guy Ron BB Moved and FPP fuse replaced
  • 240211 05/29/08 08:53 brads Powered down BB magnet
  • 240210 05/29/08 08:51 ycwang Target moved from Loop3:15cm to Loop3:6cm
  • 240207 05/29/08 08:09 ycwang Target screens
  • 240205 05/29/08 08:01 M.Olson OWL Shift Summary
  • 240201 05/29/08 07:05 M.Olson LHRS HV Screen (Run 1693)
  • 240200 05/29/08 07:04 M.Olson BigBox Screens (Run 1693)
  • 240199 05/29/08 07:03 M.Olson Beam Current Profile
  • 240198 05/29/08 07:02 M.Olson FPP Threshold and Carbon Door Screens (Run 1693)
  • 240197 05/29/08 07:01 M.Olson FPP HV and LV Screens (Run 1693)
  • 240196 05/29/08 07:00 M.Olson Hall A Tools (Run 1693)
  • 240193 05/29/08 06:42 M.Olson spot++ (Run 1692)
  • 240190 05/29/08 05:05 M.Olson Big Bite HV, restored from console
  • 240189 05/29/08 05:04 B. Hahn Target Screens
  • 240186 05/29/08 04:54 B. Hahn Target Change to 15 cm LH2
  • 240185 05/29/08 04:42 M.Olson spot++ (Run 1690)
  • 240181 05/29/08 04:13 adaq Beam Energy in A and C snapshot
  • 240180 05/29/08 04:02 adaq Target Screens snapshot
  • 240179 05/29/08 03:44 M.Olson Beam off for approximately last hour, back on now
  • 240176 05/29/08 02:12 M.Olson BB_SH map (Run 1688)
  • 240175 05/29/08 02:10 M.Olson LHRS HV (Run 1688)
  • 240174 05/29/08 02:09 M.Olson BigBox Screens (Run 1688)
  • 240173 05/29/08 02:08 M.Olson Hall A Tools (Run 1688)
  • 240171 05/29/08 01:46 B. Hahn Target back to 6 cm LH2
  • 240169 05/29/08 01:32 B. Hahn Target Moved to Dummy
  • Wednesday

  • 240163 05/28/08 23:57 Ole Hansen Swing shift summary
  • 240161 05/28/08 23:44 Ole Hansen Hall A Tools & BCM
  • 240158 05/28/08 23:08 jinhuang Target ScreenShot
  • 240156 05/28/08 23:02 Ole Hansen Beam back
  • 240154 05/28/08 23:01 Ole Hansen Brief access to reset Q1
  • 240152 05/28/08 22:27 Ole Hansen Q1 does not reset
  • 240148 05/28/08 22:24 Ole Hansen Cryo instabilities
  • 240147 05/28/08 22:21 Ole Hansen Beam back, but Q1 trip
  • 240144 05/28/08 21:59 Ole Hansen No beam - rf problems
  • 240137 05/28/08 20:13 adaq Radiation and Ozone snapshot
  • 240136 05/28/08 20:13 adaq Hall A Tools snapshot
  • 240135 05/28/08 20:13 adaq BigBox Magnet snapshot
  • 240134 05/28/08 20:13 adaq Beam Position and Current snapshot
  • 240133 05/28/08 20:13 adaq Beam Energy in A and C snapshot
  • 240132 05/28/08 20:02 adaq Target Screens snapshot
  • 240123 05/28/08 18:36 Ole Hansen VDC HV trip
  • 240120 05/28/08 17:38 jinhuang Target ScreenShot
  • 240117 05/28/08 17:33 Ole Hansen BB HV trip
  • 240114 05/28/08 17:17 xiaohui HV map K6_D2
  • 240111 05/28/08 16:30 xiaohui Acceptance check 2
  • 240110 05/28/08 16:28 xiaohui Acceptance check
  • 240109 05/28/08 16:27 xiaohui BB shower blocks K6 D2
  • 240108 05/28/08 16:24 Ole Hansen spot++ for 1st production run
  • 240107 05/28/08 16:12 jpchen Reduced Loop 3 coolant flow, increased HP limit.
  • 240105 05/28/08 16:05 Boeglin Shift Summary
  • 240094 05/28/08 14:37 Boeglin BB HV Part 2
  • 240093 05/28/08 14:36 Boeglin BB HV part 1
  • 240092 05/28/08 14:35 Boeglin Spot ++ snapshot
  • 240091 05/28/08 14:33 Boeglin BB Magnet snapshot
  • 240090 05/28/08 14:32 Boeglin BPM snapshot
  • 240089 05/28/08 14:32 Holmstrom Target Screen Capture 2
  • 240088 05/28/08 14:31 Holmstrom Target Screen 1
  • 240087 05/28/08 14:31 Boeglin Hall A Tools
  • 240086 05/28/08 14:30 Boeglin BCM snapshot
  • 240079 05/28/08 12:13 adaq Radiation and Ozone snapshot
  • 240078 05/28/08 12:13 adaq Hall A Tools snapshot
  • 240077 05/28/08 12:13 adaq BigBox Magnet snapshot
  • 240076 05/28/08 12:13 adaq Beam Position and Current snapshot
  • 240075 05/28/08 12:13 adaq Beam Energy in A and C snapshot
  • 240073 05/28/08 12:02 adaq Target Screens snapshot
  • 240070 05/28/08 11:43 Boeglin Beam quality much better now
  • 240069 05/28/08 11:25 Boeglin Beam officially back
  • 240067 05/28/08 10:30 Holmstrom Close JT valve
  • 240066 05/28/08 10:26 Boeglin No beam
  • 240064 05/28/08 10:05 Ramesh Re: T7 problem
  • 240063 05/28/08 09:42 Kalyan Follow-up on T7 problem.
  • 240062 05/28/08 09:39 Holmstrom Beam Away at 10am
  • 240051 05/28/08 08:46 Holmstrom S2 ADC 4
  • 240048 05/28/08 08:17 R. Michaels re: T7
  • 240047 05/28/08 08:02 M.Olson OWL Shift Summary
  • 240043 05/28/08 07:25 M.Olson Beam Very Unstable last 1.5 hours
  • 240040 05/28/08 06:08 M.Olson Beam Current Monitor
  • 240039 05/28/08 06:07 M.Olson FPP Low-voltage and Carbon door screens (Run 1660)
  • 240038 05/28/08 06:06 M.Olson FPP Screens (Run 1660)
  • 240037 05/28/08 06:04 M.Olson LHRS HV (Run 1660)
  • 240036 05/28/08 06:03 M.Olson BigBox Screens (Run 1660)
  • 240035 05/28/08 06:02 M.Olson Hall A Tools (Run 1660)
  • 240034 05/28/08 06:01 M.Olson spot++ (Run 1660)
  • 240031 05/28/08 05:28 M.Olson spot++ for run 1659 (Al dummy)
  • 240028 05/28/08 05:11 P. Decowski Target strip charts
  • 240025 05/28/08 04:12 adaq Radiation and Ozone snapshot
  • 240024 05/28/08 04:12 adaq Hall A Tools snapshot
  • 240023 05/28/08 04:12 adaq BigBox Magnet snapshot
  • 240022 05/28/08 04:12 adaq Beam Position and Current snapshot
  • 240021 05/28/08 04:12 adaq Beam Energy in A and C snapshot
  • 240020 05/28/08 04:02 adaq Target Screens snapshot
  • 240018 05/28/08 03:55 M.Olson BB_SH map (Kin6, 0%)
  • 240016 05/28/08 03:48 xiaohui Gas Cerenkov HV on start from 1656
  • 240015 05/28/08 03:46 xiaohui 14uA for K6 D0
  • 240014 05/28/08 03:35 xiaohui Acceptance check 2
  • 240013 05/28/08 03:33 xiaohui Acceptance check
  • 240012 05/28/08 03:30 xiaohui BB shower blocks K6 D0
  • 240006 05/28/08 02:53 Guy Ron Fuses
  • 240004 05/28/08 02:49 M.Olson spot++ for run 1653 (after access)
  • 240003 05/28/08 02:46 Guy Ron FPP Fixed
  • 240001 05/28/08 02:34 M.Olson Experts returned from access, called for beam
  • 240000 05/28/08 02:31 M.Olson BB_SH map
  • 239999 05/28/08 02:30 M.Olson LHRS HV Screen
  • 239998 05/28/08 02:30 M.Olson BigBox screens
  • 239997 05/28/08 02:28 M.Olson Hall A Tools (during access)
  • 239996 05/28/08 01:41 M.Olson Dead FPP Wire Groups Found, Called for Access, Experts Investigating
  • 239988 05/28/08 00:01 Ole Hansen Swing shift summary
  • Tuesday

  • 239987 05/27/08 23:58 Xiaohui Scint. Paddle OFF
  • 239985 05/27/08 23:46 Ole Hansen Hall A Tools & BCM log
  • 239983 05/27/08 23:35 brads Watch the data carefully!
  • 239982 05/27/08 23:16 Ole Hansen Replay results now ok
  • 239981 05/27/08 23:11 xiaohui Replay after fixing T7 (1649)
  • 239979 05/27/08 22:59 R. Michaels T7 unplugged
  • 239978 05/27/08 22:58 Ole Hansen Resuming production
  • 239972 05/27/08 22:14 Ole Hansen Trigger 7 gone wild
  • 239970 05/27/08 22:10 brads Low-E background from T7 trigger.
  • 239969 05/27/08 22:10 Ole Hansen Trigger problems
  • 239960 05/27/08 21:29 Ole Hansen Investigating BB shower ADC threshold
  • 239956 05/27/08 21:20 Ole Hansen BB Cherenkov back on
  • 239954 05/27/08 20:42 Ole Hansen BB Cherenkov off for test
  • 239950 05/27/08 20:23 xiaohui BB shower total sum ADC
  • 239949 05/27/08 20:12 adaq Radiation and Ozone snapshot
  • 239948 05/27/08 20:12 adaq Hall A Tools snapshot
  • 239947 05/27/08 20:12 adaq BigBox Magnet snapshot
  • 239946 05/27/08 20:12 adaq Beam Position and Current snapshot
  • 239945 05/27/08 20:12 adaq Beam Energy in A and C snapshot
  • 239944 05/27/08 20:02 adaq Target Screens snapshot
  • 239943 05/27/08 19:58 Ole Hansen Trippy beam
  • 239939 05/27/08 19:53 ycwang Target screens after fixed Q1 Nitrgen gas dropped
  • 239936 05/27/08 19:51 Ole Hansen Pedestal run w/beam
  • 239935 05/27/08 19:39 Ole Hansen Q1 fix
  • 239934 05/27/08 19:26 Ole Hansen Q1 problems fixed. Restoring
  • 239933 05/27/08 18:09 ycwang Q1 Nitrogen gas dropped, then the target screens
  • 239932 05/27/08 18:09 Ole Hansen Hall A Tools after Q1 trip
  • 239931 05/27/08 18:00 Ole Hansen Q1 trip
  • 239927 05/27/08 17:41 R. Michaels R-HRS Cavity status
  • 239926 05/27/08 17:41 Ole Hansen Pedestal done. Resuming production
  • 239923 05/27/08 17:39 ycwang Target moved from 6cm dummy to loop3 6cm LH2
  • 239922 05/27/08 17:35 Ole Hansen Dummy done. Start pedestal run
  • 239919 05/27/08 17:17 ycwang Target screens
  • 239918 05/27/08 17:16 ycwang Target moved from LH2 to 6cm dummy
  • 239917 05/27/08 17:11 Ole Hansen Dummy run
  • 239915 05/27/08 16:46 brads Added Cerenkov threshold readout to epicsRunStart
  • 239912 05/27/08 16:27 Ole Hansen Kin6 production in progress
  • 239908 05/27/08 16:01 Boeglin Shift Summary
  • 239904 05/27/08 15:14 Boeglin BB Magnet
  • 239903 05/27/08 15:14 Boeglin BB HV part 2
  • 239902 05/27/08 15:13 Boeglin BB HV part 1
  • 239901 05/27/08 15:09 Boeglin BPM snapshot
  • 239898 05/27/08 14:55 Holmstrom target Strip charts
  • 239897 05/27/08 14:53 Holmstrom Traget screens
  • 239896 05/27/08 14:52 Boeglin Hall A Tools snapshot and
  • 239893 05/27/08 14:49 Boeglin current raster size
  • 239891 05/27/08 14:05 Boeglin Carbon holey target run
  • 239885 05/27/08 13:00 R. Michaels R-HRS work for cavity test
  • 239881 05/27/08 12:28 V. Sulkosky Beam Energy Flat-lined
  • 239880 05/27/08 12:12 brads Controlled Access activity
  • 239879 05/27/08 12:12 adaq Radiation and Ozone snapshot
  • 239878 05/27/08 12:12 adaq Hall A Tools snapshot
  • 239877 05/27/08 12:12 adaq BigBox Magnet snapshot
  • 239876 05/27/08 12:12 adaq Beam Position and Current snapshot
  • 239875 05/27/08 12:12 adaq Beam Energy in A and C snapshot
  • 239874 05/27/08 12:02 adaq Target Screens snapshot
  • 239869 05/27/08 08:04 M.Olson OWL Shift Summary
  • 239864 05/27/08 06:48 M.Olson BB_SH Map
  • 239861 05/27/08 06:37 M.Olson spot++ (Run 1622)
  • 239860 05/27/08 06:34 M.Olson Beam Current Profile (last two hours)
  • 239859 05/27/08 06:30 xiaohui Acceptance check
  • 239858 05/27/08 06:14 xiaohui BB shower block K6 D-2
  • 239848 05/27/08 05:27 M.Olson Bigbox Screens (Run 1617)
  • 239846 05/27/08 05:24 M.Olson Beam Current Profile (Run 1617)
  • 239845 05/27/08 05:19 M.Olson BB_SH (Run 1617, Kin 6, -2%)
  • 239844 05/27/08 05:18 M.Olson Hall A Tools (Run 1617, Kin 6, -2%)
  • 239839 05/27/08 04:25 M.Olson LHRS VDC HV Trip (Run 1614)
  • 239836 05/27/08 04:12 adaq Radiation and Ozone snapshot
  • 239835 05/27/08 04:11 adaq Hall A Tools snapshot
  • 239834 05/27/08 04:11 adaq BigBox Magnet snapshot
  • 239833 05/27/08 04:11 adaq Beam Position and Current snapshot
  • 239832 05/27/08 04:11 adaq Beam Energy in A and C snapshot
  • 239831 05/27/08 04:02 adaq Target Screens snapshot
  • 239830 05/27/08 03:56 M.Olson spot++ (Run 1614)
  • 239827 05/27/08 03:27 M.Olson Target Screens
  • 239824 05/27/08 02:34 M.Olson BB HV (bbps1:2002, Run 1612)
  • 239823 05/27/08 02:33 M. Olson BB HV (bbps1:2001, Run 1612)
  • 239822 05/27/08 02:32 M.Olson BB_SH screen (Run 1612)
  • 239821 05/27/08 02:30 M.Olson LHRS HV (Run 1612)
  • 239820 05/27/08 02:29 M.Olson Beam Current Monitor (Run 1612)
  • 239819 05/27/08 02:29 M.Olson Bigbox screens (Run 1612)
  • 239818 05/27/08 02:28 M.Olson Hall A Tools (Run 1612)
  • 239817 05/27/08 02:25 M.Olson Runs 1609, 1610, 1611 Not Production
  • 239816 05/27/08 02:23 M.Olson DAQ failure at startrun of 1609
  • 239815 05/27/08 02:21 M.Olson Back to production (spot++ for run 1612)
  • 239809 05/27/08 01:10 M.Olson Secured from Moller measurement, MCC restoring beam
  • 239808 05/27/08 00:53 E.Chudakov Moller measurement
  • Monday

  • 239806 05/26/08 23:52 yez Target Screen at the end of Swing shift(2)
  • 239805 05/26/08 23:52 xqian Swing shift summary
  • 239804 05/26/08 23:51 yez Target Screen at the end of Swing shift
  • 239803 05/26/08 23:50 yez Target Loop3 Sensor_1_R Alarm
  • 239802 05/26/08 20:11 adaq Radiation and Ozone snapshot
  • 239801 05/26/08 20:11 adaq Hall A Tools snapshot
  • 239800 05/26/08 20:11 adaq BigBox Magnet snapshot
  • 239799 05/26/08 20:11 adaq Beam Position and Current snapshot
  • 239798 05/26/08 20:11 adaq Beam Energy in A and C snapshot
  • 239797 05/26/08 20:01 adaq Target Screens snapshot
  • 239796 05/26/08 19:46 yez Target changed to Empty
  • 239787 05/26/08 17:21 yez Target changed to LH2 6cm
  • 239785 05/26/08 16:51 yez Target changed to Al Dummy
  • 239782 05/26/08 16:11 yez Target Screen Swing May26(3)
  • 239781 05/26/08 16:10 yez Target Screen Swing May26(2)
  • 239780 05/26/08 16:09 yez Target Screen at the begining of Swing May25 (1)
  • 239778 05/26/08 16:03 Lindgren Shift Summary
  • 239775 05/26/08 15:50 Lindgren BigBite HV
  • 239774 05/26/08 15:49 Lindgren Energy Monitor
  • 239773 05/26/08 15:48 Lindgren Hall A Tools and current chart
  • 239772 05/26/08 15:48 Lindgren BPM/BCM
  • 239763 05/26/08 13:18 jkatich End of run failed for 1596
  • 239762 05/26/08 13:17 Lindgren End of run failed
  • 239758 05/26/08 12:11 adaq Radiation and Ozone snapshot
  • 239757 05/26/08 12:11 adaq Hall A Tools snapshot
  • 239756 05/26/08 12:11 adaq BigBox Magnet snapshot
  • 239755 05/26/08 12:11 adaq Beam Position and Current snapshot
  • 239754 05/26/08 12:11 adaq Beam Energy in A and C snapshot
  • 239753 05/26/08 12:01 adaq Target Screens snapshot
  • 239752 05/26/08 11:59 xiaohui HV map K5 D+2
  • 239751 05/26/08 11:57 Yi Qiang Errata: Production of Kin 5, delta = +2%
  • 239750 05/26/08 11:54 xiaohui Acceptance check K5_D+2
  • 239749 05/26/08 11:48 xiaohui BB shower blocks
  • 239748 05/26/08 11:37 Lindgren Production running kin 6
  • 239746 05/26/08 11:26 Lindgren Beam Current change
  • 239745 05/26/08 11:23 Lindgren Kin 6 starting set up
  • 239740 05/26/08 10:51 Lindgren Raster back on
  • 239738 05/26/08 10:43 Lindgren Spectrometer pointing
  • 239734 05/26/08 09:43 Lindgren Target change
  • 239733 05/26/08 09:42 Lindgren Production Running
  • 239731 05/26/08 09:32 Lindgren Change target
  • 239730 05/26/08 09:29 Lindgren Total Events Correction
  • 239728 05/26/08 09:16 Lindgren Day Shift Begin
  • 239724 05/26/08 08:14 M.Olson OWL Shift Summary
  • 239716 05/26/08 06:10 M.Olson BB HV (bbps1:2002, Run 1585)
  • 239715 05/26/08 06:09 M.Olson BB HV (bbps1:2001, Run 1585)
  • 239714 05/26/08 06:07 M.Olson BB_SH
  • 239713 05/26/08 06:06 M.Olson LHRS HV Screen
  • 239712 05/26/08 06:06 M.Olson Beam Current Profile (Run 1585)
  • 239711 05/26/08 06:05 M.Olson Bigbox (Run 1585)
  • 239710 05/26/08 06:04 M.Olson Hall A Tools (Run 1585)
  • 239709 05/26/08 06:03 M.Olson Spot++ (Run 1585)
  • 239704 05/26/08 04:11 adaq Radiation and Ozone snapshot
  • 239703 05/26/08 04:11 adaq Hall A Tools snapshot
  • 239702 05/26/08 04:11 adaq BigBox Magnet snapshot
  • 239701 05/26/08 04:10 adaq Beam Position and Current snapshot
  • 239700 05/26/08 04:10 adaq Beam Energy in A and C snapshot
  • 239697 05/26/08 04:01 adaq Target Screens snapshot
  • 239690 05/26/08 01:18 M.Olson spot++ (Run 1579)
  • 239689 05/26/08 01:16 M.Olson Target Screens (Run 1579)
  • 239688 05/26/08 01:14 M.Olson LHRS HV Screen (run 1579)
  • 239687 05/26/08 01:13 M.Olson Bigbox screen (run 1579)
  • 239686 05/26/08 01:13 M.Olson Beam Current Monitor
  • 239685 05/26/08 01:12 M.Olson Hall A Tools
  • 239684 05/26/08 01:10 M.Olson BB HV Screens (bbps2:2002)
  • 239683 05/26/08 01:09 M.Olson BB HV Screens (bbps1:2001)
  • 239682 05/26/08 01:08 M.Olson BB_SH
  • 239678 05/26/08 00:09 Ramesh Shift Summary
  • 239676 05/26/08 00:07 Ramesh ROC5 reboot
  • Sunday

  • 239673 05/25/08 22:22 xiaohui 18uA for K5-D0 start from 1576
  • 239671 05/25/08 22:07 xiaohui BB shower block hit pattern K4 D0
  • 239668 05/25/08 21:38 xiaohui BigBite shower block tests
  • 239664 05/25/08 21:21 brads Re-enabled the Cerenkov PMTs, updated HV gui files
  • 239662 05/25/08 20:51 Ramesh No beam for a while without a known reason
  • 239656 05/25/08 20:10 adaq Radiation and Ozone snapshot
  • 239655 05/25/08 20:10 adaq Hall A Tools snapshot
  • 239654 05/25/08 20:10 adaq BigBox Magnet snapshot
  • 239653 05/25/08 20:10 adaq Beam Position and Current snapshot
  • 239652 05/25/08 20:10 adaq Beam Energy in A and C snapshot
  • 239650 05/25/08 20:01 adaq Target Screens snapshot
  • 239648 05/25/08 19:57 Ramesh Ready for delta=0% at Kin 5
  • 239639 05/25/08 19:04 Ramesh Beam trip level very high
  • 239636 05/25/08 17:54 Kalyan Target Screens
  • 239634 05/25/08 17:50 Ramesh BigBite HV tripped
  • 239632 05/25/08 17:26 Ramesh Tech: Q2 liquid level redback red
  • 239629 05/25/08 16:49 Ramesh spot check
  • 239627 05/25/08 16:33 Lindgren Shift Summary
  • 239623 05/25/08 15:05 Lindgren Dummy Target
  • 239618 05/25/08 13:47 Xiaohui Acceptance check 2
  • 239617 05/25/08 13:46 Xiaohui Acceptance check for K4_D-2
  • 239616 05/25/08 13:22 Lindgren Beam Energy
  • 239615 05/25/08 13:21 Lindgren BB_SH
  • 239614 05/25/08 13:20 Lindgren BPM/BCM
  • 239613 05/25/08 13:18 Lindgren BigBox
  • 239612 05/25/08 13:18 Lindgren Hall A Tools/Current monitor
  • 239611 05/25/08 13:14 Lindgren FSD Trips
  • 239608 05/25/08 12:37 Norum Target status
  • 239607 05/25/08 12:33 Norum Target status
  • 239606 05/25/08 12:10 adaq Radiation and Ozone snapshot
  • 239605 05/25/08 12:10 adaq Hall A Tools snapshot
  • 239604 05/25/08 12:10 adaq BigBox Magnet snapshot
  • 239603 05/25/08 12:10 adaq Beam Position and Current snapshot
  • 239602 05/25/08 12:10 adaq Beam Energy in A and C snapshot
  • 239601 05/25/08 12:01 adaq Target Screens snapshot
  • 239596 05/25/08 10:08 xiaohui FPP HV monitor
  • 239595 05/25/08 10:07 Xiaohui Increase FPP HV to 1850
  • 239590 05/25/08 08:46 Lindgren Day shift begins
  • 239589 05/25/08 08:43 Lindgren Day shift begins
  • 239585 05/25/08 08:01 Reinhold shift summary owl shift
  • 239582 05/25/08 07:46 Guy Ron BigBite HV Shower map for K5 D=-2%
  • 239581 05/25/08 07:43 Guy Ron BigBite elastic hit position
  • 239580 05/25/08 07:41 Reinhold Kinematic 5: -2% set
  • 239578 05/25/08 07:30 reinhold magnets are settling; ask for beam back
  • 239577 05/25/08 07:17 Reinhold finished access: moved big bite to 40 degree
  • 239576 05/25/08 06:57 M.Olson Target moved to LH2 (6cm)
  • 239574 05/25/08 06:54 Reinhold Take beam away for kinematic change
  • 239572 05/25/08 06:34 M.Olson Target moved to Al Dummy (6 cm)
  • 239571 05/25/08 06:31 Reinhold Finished production for Kin 4 2% with 68M
  • 239569 05/25/08 06:04 reinhold plan for change to next kinematic at ~ 7am
  • 239566 05/25/08 05:40 M.Olson Target Screens
  • 239563 05/25/08 05:07 M Olson spot++ for run 1547
  • 239560 05/25/08 04:10 adaq Radiation and Ozone snapshot
  • 239559 05/25/08 04:10 adaq Hall A Tools snapshot
  • 239558 05/25/08 04:10 adaq BigBox Magnet snapshot
  • 239557 05/25/08 04:10 adaq Beam Position and Current snapshot
  • 239556 05/25/08 04:09 adaq Beam Energy in A and C snapshot
  • 239555 05/25/08 04:01 adaq Target Screens snapshot
  • 239544 05/25/08 01:02 brads Beam energy readback flat-lined since 8pm
  • 239541 05/25/08 00:12 reinhold adjusted ps4 for run 1541
  • Saturday

  • 239537 05/24/08 23:55 Ramesh Swing Shift Summary
  • 239536 05/24/08 23:52 Y. Zhang target screen.
  • 239533 05/24/08 22:53 Y. Zhang B.B. HV tripped
  • 239528 05/24/08 21:47 xiaohui Target Y & coin. time
  • 239527 05/24/08 21:45 xiaohui Acceptance check
  • 239526 05/24/08 21:39 Xiaohui BB HV map K4-D2
  • 239525 05/24/08 21:38 xiaohui BB shower blocks K4-D2
  • 239514 05/24/08 20:09 adaq Radiation and Ozone snapshot
  • 239513 05/24/08 20:09 adaq Hall A Tools snapshot
  • 239512 05/24/08 20:09 adaq BigBox Magnet snapshot
  • 239511 05/24/08 20:09 adaq Beam Position and Current snapshot
  • 239510 05/24/08 20:09 adaq Beam Energy in A and C snapshot
  • 239509 05/24/08 20:08 Ramesh spot check during shift check-list
  • 239508 05/24/08 20:01 adaq Target Screens snapshot
  • 239505 05/24/08 19:17 brads BB Cerenkov HV settings
  • 239494 05/24/08 16:35 Ramesh DAQ reboot (late entry)
  • 239492 05/24/08 16:05 Lindgren Day Shift Summary
  • 239489 05/24/08 14:56 Lindgren Spot++
  • 239488 05/24/08 14:52 Lindgren BB_SH
  • 239487 05/24/08 14:51 Lindgren BB_SH
  • 239486 05/24/08 14:50 Lindgren BPM
  • 239485 05/24/08 14:47 Lindgren Hall A Tools and Current Monitor
  • 239484 05/24/08 14:46 Norum Target status
  • 239483 05/24/08 14:45 Norum Target status
  • 239474 05/24/08 12:11 xiaohui Procedure for selecting shower blocks
  • 239473 05/24/08 12:09 adaq Radiation and Ozone snapshot
  • 239472 05/24/08 12:09 adaq Hall A Tools snapshot
  • 239471 05/24/08 12:09 adaq BigBox Magnet snapshot
  • 239470 05/24/08 12:09 adaq Beam Position and Current snapshot
  • 239469 05/24/08 12:09 adaq Beam Energy in A and C snapshot
  • 239468 05/24/08 12:01 adaq Target Screens snapshot
  • 239465 05/24/08 11:47 Yi Qiang Screen of Active Shower Blocks
  • 239464 05/24/08 11:45 Yi Qiang Active Shower Block Map
  • 239463 05/24/08 11:43 Xiaohui Coin. Time
  • 239462 05/24/08 11:33 xiaohui LHRS Acceptance for 6 blocks and 8 blocks
  • 239461 05/24/08 11:24 Lindgren Shower Counter
  • 239459 05/24/08 11:06 R.Lindgren Shower Counter
  • 239456 05/24/08 10:50 R. Lindgren Shower counter
  • 239454 05/24/08 10:38 R. Lindgren change to delta=0% Kinematics
  • 239450 05/24/08 09:59 Guy Ron BB HV Map K4 D0
  • 239449 05/24/08 09:57 Guy Ron BB Hit Map K4 D=0
  • 239445 05/24/08 09:24 Lindgren Target change
  • 239443 05/24/08 09:13 R. Lindgren Change target
  • 239437 05/24/08 08:11 Norum Target status
  • 239436 05/24/08 08:06 Reinhold owl shift summary
  • 239430 05/24/08 06:34 Reinhold Hall A Tools
  • 239427 05/24/08 06:23 Reinhold SPOT for shift checklist
  • 239424 05/24/08 05:30 A. Camsonne Loop3 fan current settings
  • 239423 05/24/08 05:16 A. Camsonne Fan frequency fluctuation
  • 239421 05/24/08 05:10 reinhold FFB tripped
  • 239417 05/24/08 04:09 adaq Radiation and Ozone snapshot
  • 239416 05/24/08 04:09 adaq Hall A Tools snapshot
  • 239415 05/24/08 04:09 adaq BigBox Magnet snapshot
  • 239414 05/24/08 04:09 adaq Beam Position and Current snapshot
  • 239413 05/24/08 04:09 adaq Beam Energy in A and C snapshot
  • 239412 05/24/08 04:01 adaq Target Screens snapshot
  • 239408 05/24/08 03:11 A. Camsonne Error message when rebooting TS12
  • 239407 05/24/08 03:05 A. Camsonne EB warning message for run 1502
  • 239405 05/24/08 03:03 A. Camsonne Run 1502 100% deadtime
  • 239404 05/24/08 02:33 A. Camsonne Temperature / beam excursion
  • 239403 05/24/08 02:32 reinhold run 1501: beam position unstable
  • 239400 05/24/08 02:26 Reinhold FFB reboot done; beam back
  • 239399 05/24/08 02:20 Reinhold Run 1500: FBB reboot; probably junk
  • 239396 05/24/08 02:17 reinhold let MCC reboot FFB (fast feedback) IOC
  • 239393 05/24/08 01:42 reinhold MCC calls about feedback not working
  • 239390 05/24/08 01:12 A. Camsonne Target screens
  • 239389 05/24/08 01:03 A. Camsonne Target screens
  • 239388 05/24/08 01:03 A. Camsonne Target moved to 6cm LH2
  • 239386 05/24/08 00:52 reinhold spot check with holey carbon
  • 239383 05/24/08 00:47 Reinhold Beam back, running C w/hole
  • 239378 05/24/08 00:31 A. Camsonne Target moved to carbon hole
  • 239377 05/24/08 00:02 Boeglin Shift Summary
  • Friday

  • 239368 05/23/08 20:09 adaq Radiation and Ozone snapshot
  • 239367 05/23/08 20:08 adaq Hall A Tools snapshot
  • 239366 05/23/08 20:08 adaq BigBox Magnet snapshot
  • 239365 05/23/08 20:08 adaq Beam Position and Current snapshot
  • 239364 05/23/08 20:08 adaq Beam Energy in A and C snapshot
  • 239362 05/23/08 20:01 adaq Target Screens snapshot
  • 239360 05/23/08 19:42 Yi Qiang Repost Acceptance plot with a visible title
  • 239359 05/23/08 19:40 Xiaohui HV map K4_D0
  • 239358 05/23/08 19:38 Yi Qiang Acceptance Check of Run 1490
  • 239357 05/23/08 19:34 Xiaohui BB shower blocks K4_D-2
  • 239356 05/23/08 19:26 Yi Qiang Hall A Tools Before Kin4
  • 239354 05/23/08 19:20 Yi Qiang Spot++ of run 1489
  • 239348 05/23/08 18:21 Norum Target status
  • 239338 05/23/08 16:01 Saha Shift Summary
  • 239333 05/23/08 15:00 hyao target charts
  • 239323 05/23/08 12:08 adaq Radiation and Ozone snapshot
  • 239322 05/23/08 12:08 adaq Hall A Tools snapshot
  • 239321 05/23/08 12:08 adaq BigBox Magnet snapshot
  • 239320 05/23/08 12:08 adaq Beam Position and Current snapshot
  • 239319 05/23/08 12:08 adaq Beam Energy in A and C snapshot
  • 239318 05/23/08 12:01 adaq Target Screens snapshot
  • 239311 05/23/08 09:32 Yi Qiang BB Shower HV Trip
  • 239308 05/23/08 08:59 Yi Qiang Acceptance of Run 1472
  • 239304 05/23/08 08:08 S. Beck 23-May Owl shift summary
  • 239300 05/23/08 07:49 Guy Ron BB HV Map K3 D=+2
  • 239295 05/23/08 07:23 S. Beck Efficiency problem solved
  • 239294 05/23/08 07:20 S. Beck Beam spot problem solved
  • 239290 05/23/08 06:38 S. Beck The beam spot is spread to 9x9 !!!
  • 239288 05/23/08 06:10 adaq A problem - efficiency drop in BB and FPP WCs
  • 239285 05/23/08 05:22 S. Beck Monitoring the FPP WC efficiencies
  • 239284 05/23/08 05:17 Chiranjib Shower blocks check
  • 239281 05/23/08 04:11 Chiranjib Raster
  • 239280 05/23/08 04:08 adaq Radiation and Ozone snapshot
  • 239279 05/23/08 04:08 adaq Hall A Tools snapshot
  • 239278 05/23/08 04:08 adaq BigBox Magnet snapshot
  • 239277 05/23/08 04:08 adaq Beam Position and Current snapshot
  • 239276 05/23/08 04:08 adaq Beam Energy in A and C snapshot
  • 239273 05/23/08 04:00 adaq Target Screens snapshot
  • 239272 05/23/08 03:53 S. Beck Misleading information in "FPP Carbon Doors" window
  • 239271 05/23/08 03:26 jamesj Deadtime window disappeared, brought back
  • 239268 05/23/08 03:11 Chiranjib TS11 rebooted
  • 239265 05/23/08 02:43 S. Beck, Magnet mismatch on BB
  • 239264 05/23/08 02:11 S. Beck VDCs HV tripped again.
  • 239258 05/23/08 00:17 Chiranjib TARGET SCREENS
  • 239257 05/23/08 00:02 LeRose Shift Summary (swing)
  • Thursday

  • 239242 05/22/08 21:08 Reinhold Moved target from 6 cm dummy to 6cm LH2 (loop 3)
  • 239240 05/22/08 20:57 LeRose VDC HV trip
  • 239236 05/22/08 20:33 Reinhold Moved target from 6 cm LH2 to 6cm dummy
  • 239234 05/22/08 20:08 adaq Radiation and Ozone snapshot
  • 239233 05/22/08 20:08 adaq Hall A Tools snapshot
  • 239232 05/22/08 20:07 adaq BigBox Magnet snapshot
  • 239231 05/22/08 20:07 adaq Beam Position and Current snapshot
  • 239230 05/22/08 20:07 adaq Beam Energy in A and C snapshot
  • 239229 05/22/08 20:00 adaq Target Screens snapshot
  • 239224 05/22/08 18:58 Xiaohui HV map
  • 239223 05/22/08 18:55 xiaohui Fpp efficiency
  • 239222 05/22/08 18:53 Xiaohui BigBite Shower Blocks at K3_D0
  • 239216 05/22/08 18:13 Reinhold Moved target from tantalum to 6 cm LH2 (hydrogen) in Loop 3
  • 239213 05/22/08 18:05 Reinhold Moved target from carbon to tantalum
  • 239210 05/22/08 17:52 Reinhold Moved target from carbon w/hole to carbon target
  • 239208 05/22/08 17:49 LeRose Spot++ Holey Carbon
  • 239204 05/22/08 17:34 LeRose Beam Back
  • 239203 05/22/08 16:09 Reinhold Target screen shots #2
  • 239202 05/22/08 16:07 Reinhold Target screen shots #1
  • 239200 05/22/08 16:01 Saha Shift Summary
  • 239199 05/22/08 15:40 spiegel/Tech Beam line vacuum valves failure
  • 239198 05/22/08 14:32 Xiaohui FPP recovered
  • 239197 05/22/08 14:23 hyao Taget moved from H2 to Carbon(with hole)
  • 239196 05/22/08 13:09 adaq beam back on
  • 239195 05/22/08 12:07 adaq Radiation and Ozone snapshot
  • 239194 05/22/08 12:07 adaq Hall A Tools snapshot
  • 239193 05/22/08 12:07 adaq BigBox Magnet snapshot
  • 239192 05/22/08 12:07 adaq Beam Position and Current snapshot
  • 239191 05/22/08 12:07 adaq Beam Energy in A and C snapshot
  • 239190 05/22/08 12:00 adaq Target Screens snapshot
  • 239189 05/22/08 11:18 Guy Ron FPP LV "problem" + Carbon doors
  • 239187 05/22/08 08:02 E.Chudakov Owl shift summary
  • 239185 05/22/08 07:28 E.Chudakov FPP LV power, screen
  • 239182 05/22/08 07:18 E.Chudakov Staying with the same settings
  • 239181 05/22/08 07:01 E.Chudakov Moller results for Wien angle -33 and -13 degrees
  • 239177 05/22/08 05:38 S. Beck A problem with two FPP 5V low voltage
  • 239171 05/22/08 04:09 E.Chudakov MCC - cavity problems
  • 239170 05/22/08 04:07 adaq Radiation and Ozone snapshot
  • 239169 05/22/08 04:07 adaq Hall A Tools snapshot
  • 239168 05/22/08 04:07 adaq BigBox Magnet snapshot
  • 239167 05/22/08 04:07 adaq Beam Position and Current snapshot
  • 239166 05/22/08 04:07 adaq Beam Energy in A and C snapshot
  • 239165 05/22/08 04:00 adaq Target Screens snapshot
  • 239164 05/22/08 03:58 S. Beck VDCs V1 and V2 look a bit different
  • 239161 05/22/08 03:45 S. Beck 6th scintillator of S1,S2 don't see events
  • 239159 05/22/08 02:52 E.Chudakov BB HV on
  • 239158 05/22/08 02:48 jamesj Target: moved to LH2
  • 239157 05/22/08 02:31 E.Chudakov Moller measurement finished
  • 239156 05/22/08 01:14 E.Chudakov Moller data taking started
  • 239155 05/22/08 00:39 E.Chudakov BB HV turned off
  • 239154 05/22/08 00:31 E.Chudakov Starting Moller measurement
  • 239153 05/22/08 00:23 jamesj Target: moved to empty target
  • Wednesday

  • 239150 05/21/08 23:58 Seonho Choi Shift Summary (Swing)
  • 239147 05/21/08 23:23 Xiaohui FPP 0.75 inch door not closed
  • 239145 05/21/08 23:12 Reinhold Moved target from 6 cm dummy to 6cm LH2 (loop 3)
  • 239143 05/21/08 22:45 Seonho Choi Hall-A Tools Screen Capture
  • 239141 05/21/08 22:41 reinhold Moved target from 6cm LH2 to 6cm dummy
  • 239135 05/21/08 21:06 xiaohui Axis name correction
  • 239134 05/21/08 21:02 Xiaohui FPP efficiency is coming back
  • 239133 05/21/08 20:48 brads Big beam energy drift in last hour
  • 239130 05/21/08 20:07 adaq Radiation and Ozone snapshot
  • 239129 05/21/08 20:07 adaq Hall A Tools snapshot
  • 239128 05/21/08 20:07 adaq BigBox Magnet snapshot
  • 239127 05/21/08 20:06 adaq Beam Position and Current snapshot
  • 239126 05/21/08 20:06 adaq Beam Energy in A and C snapshot
  • 239125 05/21/08 20:00 adaq Target Screens snapshot
  • 239119 05/21/08 18:32 Seonho Choi Incorrect label for run 1423
  • 239113 05/21/08 16:58 adaq target screen shots #2
  • 239112 05/21/08 16:57 adaq target screen shot #1
  • 239108 05/21/08 16:06 sarty Day Shift Summary
  • 239107 05/21/08 16:03 R. Michaels Run 1419 to check FPP
  • 239102 05/21/08 15:52 R. Michaels test runs 1417, 1418
  • 239097 05/21/08 15:26 Xiaohui FPP efficiency keep dropping
  • 239096 05/21/08 14:29 brads BigBite now at 35 degrees
  • 239095 05/21/08 14:14 sarty MCC trying to track down causes of BLM trips
  • 239094 05/21/08 14:08 R. Michaels opened JT valve 3 clicks
  • 239093 05/21/08 13:50 R. Michaels Run 1416 with restricted BB trig.
  • 239091 05/21/08 13:47 Guy Ron BB High Voltage map
  • 239090 05/21/08 13:45 sarty Running with no "fast feedback" for beam
  • 239089 05/21/08 13:43 R. Michaels Run 1415 (kine 3)
  • 239087 05/21/08 13:40 Jack (Techs) Gas system low pressure alarm
  • 239085 05/21/08 13:36 sarty HRS-L angle now set for Kin3
  • 239084 05/21/08 13:21 sarty Set HRS-L and BigBite for Kin3 (delta=-2%)
  • 239083 05/21/08 13:12 R. Michaels R-HRS DAQ revived (4 cavity test)
  • 239082 05/21/08 13:00 brads BB ramped to 0 for check, ramped back to 700A now
  • 239080 05/21/08 12:52 Guy Ron BigBite move completed
  • 239079 05/21/08 12:11 sarty Finished Kin2 (all delta settings)
  • 239078 05/21/08 12:06 adaq Radiation and Ozone snapshot
  • 239077 05/21/08 12:06 adaq Hall A Tools snapshot
  • 239076 05/21/08 12:06 adaq BigBox Magnet snapshot
  • 239075 05/21/08 12:06 adaq Beam Position and Current snapshot
  • 239074 05/21/08 12:06 adaq Beam Energy in A and C snapshot
  • 239072 05/21/08 12:00 adaq Target Screens snapshot
  • 239070 05/21/08 11:46 Jack (Techs) Nitrogen line not charged
  • 239068 05/21/08 11:33 R. Michaels gas restored
  • 239067 05/21/08 11:24 sarty Gas Shed problems again
  • 239066 05/21/08 11:20 Jack (Techs) Gas System
  • 239063 05/21/08 10:37 Guy Ron FPP Gas Problems
  • 239062 05/21/08 10:37 Guy Ron FPP Gas Problems
  • 239059 05/21/08 10:10 R. Michaels gas is flowing
  • 239058 05/21/08 10:04 R. Michaels Gas shed power failure / solution /implication
  • 239052 05/21/08 08:39 R. Michaels updated DAQ documentation.
  • 239050 05/21/08 08:19 S. Beck 5/21/08 owl shift summary
  • 239042 05/21/08 06:27 S. Beck X/Y positions on BPMA were changed
  • 239039 05/21/08 05:29 S. Beck raster size is too big
  • 239034 05/21/08 04:06 adaq Radiation and Ozone snapshot
  • 239033 05/21/08 04:06 adaq Hall A Tools snapshot
  • 239032 05/21/08 04:06 adaq BigBox Magnet snapshot
  • 239031 05/21/08 04:06 adaq Beam Position and Current snapshot
  • 239030 05/21/08 04:06 adaq Beam Energy in A and C snapshot
  • 239029 05/21/08 04:00 adaq Target Screens snapshot
  • 239026 05/21/08 03:42 S. Beck No information about Gas flow (GAS SHED)
  • 239023 05/21/08 03:03 Xiaohui Reboot ROC 5
  • 239022 05/21/08 02:55 S. Beck Problems with CODA caused by ROC5
  • 239019 05/21/08 01:25 ycwang cannot ending a run
  • 239015 05/21/08 00:35 ycwang 1394 is junk
  • Tuesday

  • 239009 05/20/08 23:55 Seonho Choi Shift Summary (Swing)
  • 239008 05/20/08 23:54 Seonho Choi Beam instability. MCC is investigating.
  • 239007 05/20/08 23:38 Seonho Choi Hall A Tools Capture
  • 239006 05/20/08 23:37 Seonho Choi Raster from the run just after recovery
  • 239004 05/20/08 23:32 Seonho Choi Beam On Target
  • 239002 05/20/08 23:12 Seonho Choi Magnet Mismatch from MCC
  • 239001 05/20/08 23:10 Michaels / G daq problems / repair
  • 239000 05/20/08 23:05 Seonho Choi TS11 fixed finally
  • 238997 05/20/08 22:06 Xiaohui BB shower hits
  • 238996 05/20/08 21:57 Yi Qiang Controlled Access to reboot TS11
  • 238995 05/20/08 21:55 Seonho Choi Power Supply Problem: Bill Gunning has been contacted, to be here by 22:30
  • 238994 05/20/08 21:40 Seonho Choi Power supply problem (cont)
  • 238993 05/20/08 21:36 Seonho Choi Power supply for TS11 mal-functioning
  • 238992 05/20/08 20:51 Seonho Choi Controlled Access to reset TS11
  • 238991 05/20/08 20:35 Seonho Choi TS11 Crash, recovery in progress
  • 238990 05/20/08 20:25 Chiranjib Raster check
  • 238988 05/20/08 20:13 chiranjib TARGET--->> L3
  • 238987 05/20/08 20:08 Yi Qiang Carbon Hole spot++
  • 238986 05/20/08 20:06 adaq Radiation and Ozone snapshot
  • 238985 05/20/08 20:06 adaq Hall A Tools snapshot
  • 238984 05/20/08 20:06 adaq BigBox Magnet snapshot
  • 238983 05/20/08 20:06 adaq Beam Position and Current snapshot
  • 238982 05/20/08 20:06 adaq Beam Energy in A and C snapshot
  • 238978 05/20/08 20:00 adaq Target Screens snapshot
  • 238977 05/20/08 19:58 brads Brief Controlled Access to look at BB Cerenkov gas system
  • 238973 05/20/08 18:11 xiaohui Energy Deposition in PS
  • 238972 05/20/08 17:53 Kalyan Thresholds on Bigbite scintillator plane
  • 238971 05/20/08 17:47 Chiranjib TARGET ----> C-hole
  • 238970 05/20/08 16:02 sarty Shift Summary
  • 238968 05/20/08 14:36 R. Michaels request about R-HRS
  • 238967 05/20/08 14:17 Xiaohui Database changed db_DL.dat
  • 238966 05/20/08 14:12 sarty HRS-L Quads all tripped off - going back down to Reset
  • 238965 05/20/08 13:57 Xiaohui Clarification on TDC variables
  • 238963 05/20/08 13:22 Moffit ROC5 added to TWOARM_FB
  • 238949 05/20/08 12:05 adaq Radiation and Ozone snapshot
  • 238948 05/20/08 12:05 adaq Hall A Tools snapshot
  • 238947 05/20/08 12:05 adaq BigBox Magnet snapshot
  • 238946 05/20/08 12:05 adaq Beam Position and Current snapshot
  • 238945 05/20/08 12:05 adaq Beam Energy in A and C snapshot
  • 238943 05/20/08 12:00 adaq Target Screens snapshot
  • 238942 05/20/08 11:54 hyao At least 1 more hour(MCC update)
  • 238937 05/20/08 10:51 hyao turn HV 2002 Off
  • 238936 05/20/08 10:26 sarty Machine down (since ~09:00) due to Thunderstorm knocking out PSS and Accel. Power Suppiles
  • 238935 05/20/08 10:06 hyao Controlled Access(Now)
  • 238934 05/20/08 08:43 hyao Target moved from 6cm dummy to Loop3 6cm H2
  • 238933 05/20/08 08:03 E.Chudakov Owl shift summary
  • 238931 05/20/08 07:43 E.Chudakov Update from MCC
  • 238929 05/20/08 04:05 adaq Radiation and Ozone snapshot
  • 238928 05/20/08 04:05 adaq Hall A Tools snapshot
  • 238927 05/20/08 04:05 adaq BigBox Magnet snapshot
  • 238926 05/20/08 04:05 adaq Beam Position and Current snapshot
  • 238925 05/20/08 04:05 adaq Beam Energy in A and C snapshot
  • 238924 05/20/08 04:00 adaq Target Screens snapshot
  • 238923 05/20/08 03:47 E.Chudakov No beam seen for 1h
  • 238919 05/20/08 02:24 Armstrong Target: move to Al dummy
  • 238917 05/20/08 02:19 Armstrong replay: noisy FPP channels
  • 238916 05/20/08 02:19 Xiaohui Important Info about Wierd Coin. Time Shape
  • 238913 05/20/08 01:44 xiaohui BB shower block 1.7(adc 161)
  • 238910 05/20/08 01:03 Armstrong Target: screen shots
  • 238909 05/20/08 01:03 xiaohui Plots
  • 238908 05/20/08 00:58 xiaohui BB shower threshold scan
  • 238907 05/20/08 00:57 xiaohui BB shower threshold scan
  • 238906 05/20/08 00:52 brads Updated halog command sucks less
  • 238903 05/20/08 00:42 brads Coin. time mysteries (repost)
  • 238895 05/20/08 00:02 sarty Swing Shift Summary
  • Monday

  • 238891 05/19/08 23:34 xiaohui Decrease threshold
  • 238889 05/19/08 23:27 xiaohui About the drop in coint.time spectrum
  • 238888 05/19/08 22:52 brads Coin. time mysteries and musings
  • 238887 05/19/08 22:50 brads Turning off PSh blocks == increasing effective disc. threshold for BB trigger
  • 238880 05/19/08 21:27 sarty Optimizing which BB shower blocks to keep on (runs 1359-1361)
  • 238879 05/19/08 21:18 Nikos effect of Preshowers on coincidence time
  • 238868 05/19/08 20:05 adaq Radiation and Ozone snapshot
  • 238867 05/19/08 20:05 adaq Hall A Tools snapshot
  • 238866 05/19/08 20:05 adaq BigBox Magnet snapshot
  • 238865 05/19/08 20:05 adaq Beam Position and Current snapshot
  • 238864 05/19/08 20:05 adaq Beam Energy in A and C snapshot
  • 238863 05/19/08 20:00 adaq Target Screens snapshot
  • 238862 05/19/08 19:51 Nikos accidentals
  • 238858 05/19/08 18:57 Xiaohui Turn on all BB shower blocks starting 1357
  • 238857 05/19/08 18:42 Nikos accidentals in coincidence time
  • 238855 05/19/08 18:15 sarty Started Kin2, delta= 0% production (BB Wire Chamber studies completed)
  • 238837 05/19/08 17:42 xqian HV 1750 V 6 cm LH2 1uA
  • 238828 05/19/08 17:37 xqian HV 1700 V 1uA 6cm LH2
  • 238827 05/19/08 17:36 ycwang Target Moved back to LH2
  • 238819 05/19/08 17:33 xqian HV 1650 V 0.98 uA LH2 6cm
  • 238809 05/19/08 17:27 xqian HV current 0.85 uA
  • 238808 05/19/08 17:23 R. Michaels error -- no sign check done
  • 238805 05/19/08 16:50 ycwang Target moved from LH2 to 6cm Dummy
  • 238803 05/19/08 16:39 Huber Run 1333 - FPP WG on U1
  • 238802 05/19/08 16:15 ycwang Target screens
  • 238801 05/19/08 16:14 ycwang Target Screens
  • 238797 05/19/08 16:02 jpchen Day shift summary
  • 238796 05/19/08 15:55 Kalyan Bigbite calorimeter plots - run 1285
  • 238795 05/19/08 15:41 R. Michaels mini-sign check.
  • 238794 05/19/08 15:33 jpchen Left Collimator moved to OPEN
  • 238792 05/19/08 15:26 R. Michaels opening collimator now
  • 238790 05/19/08 15:24 Xiaohui BB shower block 1.7 was on during 1325,1326
  • 238789 05/19/08 15:14 Kalyan Some BB calorimeter plots
  • 238788 05/19/08 15:07 R. Michaels scaler checks, Qasy check
  • 238787 05/19/08 14:43 R. Michaels L-HRS scint ADC high
  • 238786 05/19/08 14:33 R. Michaels resume production
  • 238785 05/19/08 14:33 jpchen BigBite back to 700 A
  • 238783 05/19/08 14:23 jpchen BigBite Magnet field to 0
  • 238782 05/19/08 14:23 R. Michaels stop run 1331 -- do Moller next
  • 238780 05/19/08 14:13 jpchen LHRS collimator is in 6 msr position
  • 238779 05/19/08 13:56 eschulte Xin Qian's BigBite Wire Chamber Test for Monday 19 May
  • 238776 05/19/08 12:56 S.Beck strange feature in BB sum TDC hist.
  • 238775 05/19/08 12:50 R. Michaels screen captures of target
  • 238774 05/19/08 12:49 R. Michaels wrong start-run info (target)
  • 238771 05/19/08 12:41 S. Beck Bad L-arm S2 ADDC#5 ?
  • 238770 05/19/08 12:17 S. Beck LHRS scintillator ADCs saturating?
  • 238769 05/19/08 12:05 adaq Radiation and Ozone snapshot
  • 238768 05/19/08 12:05 adaq Hall A Tools snapshot
  • 238767 05/19/08 12:04 adaq BigBox Magnet snapshot
  • 238766 05/19/08 12:04 adaq Beam Position and Current snapshot
  • 238765 05/19/08 12:04 adaq Beam Energy in A and C snapshot
  • 238762 05/19/08 11:59 adaq Target Screens snapshot
  • 238761 05/19/08 11:46 R. Michaels run number anomalies e04007
  • 238758 05/19/08 10:43 R. Michaels missing replay info
  • 238757 05/19/08 10:41 R. Michaels origin of neg. DT w/beam off
  • 238752 05/19/08 10:23 R. Michaels Peaks in BB Tsum ADC
  • 238751 05/19/08 10:13 R. Michaels run 1324 is production
  • 238748 05/19/08 09:53 R. Michaels deadtime decreased 6% (pedsup ROC8)
  • 238731 05/19/08 08:01 Solvignon Shift summary
  • 238725 05/19/08 06:43 Armstrong need official golden run
  • 238724 05/19/08 06:37 Armstrong Target: possible diagnosis
  • 238723 05/19/08 06:06 Armstrong Target: brief alarms
  • 238720 05/19/08 05:04 Guy Ron BB after Kinematic change
  • 238719 05/19/08 05:03 Guy Ron BB after Kinematic change
  • 238716 05/19/08 04:49 Solvignon Kin2 delta=-2% start
  • 238713 05/19/08 04:04 adaq Radiation and Ozone snapshot
  • 238712 05/19/08 04:04 adaq Hall A Tools snapshot
  • 238711 05/19/08 04:04 adaq BigBox Magnet snapshot
  • 238710 05/19/08 04:04 adaq Beam Position and Current snapshot
  • 238709 05/19/08 04:04 adaq Beam Energy in A and C snapshot
  • 238708 05/19/08 03:59 adaq Target Screens snapshot
  • 238707 05/19/08 03:49 Solvignon Spot++
  • 238706 05/19/08 03:45 Armstrong Taregt: screen shots
  • 238703 05/19/08 03:11 xiaohui Carbon run 1290 with DT 99%
  • 238696 05/19/08 00:20 Armstrong Target status
  • 238693 05/19/08 00:03 eschulte Beam Trip Issues
  • Sunday

  • 238691 05/18/08 23:58 Ramesh Subed Shift Summary
  • 238690 05/18/08 23:13 Ramesh spot++ during shift check-list
  • 238689 05/18/08 23:08 Ramesh Beam back
  • 238687 05/18/08 23:00 Ramesh Beam is still off
  • 238686 05/18/08 22:59 Kalyan Target screens
  • 238685 05/18/08 22:07 Ramesh Beam problem persists
  • 238684 05/18/08 21:46 Ramesh Brief beam
  • 238681 05/18/08 21:18 jpchen Power glitch cause fan sop. Reset fine now.
  • 238680 05/18/08 21:15 Ramesh Target ready, ready to take beam
  • 238679 05/18/08 20:50 Ramesh Asked for controlled access
  • 238678 05/18/08 20:37 Kalyan Loop 3 fan tripped
  • 238677 05/18/08 20:04 adaq Radiation and Ozone snapshot
  • 238676 05/18/08 20:04 adaq Hall A Tools snapshot
  • 238675 05/18/08 20:04 adaq BigBox Magnet snapshot
  • 238674 05/18/08 20:04 adaq Beam Position and Current snapshot
  • 238673 05/18/08 20:04 adaq Beam Energy in A and C snapshot
  • 238672 05/18/08 20:01 Ramesh Beam stops
  • 238666 05/18/08 19:28 Ramesh Trippy beam
  • 238661 05/18/08 17:56 xiaohui clarification about coint. time cut
  • 238660 05/18/08 17:42 xiaohui Elastic vs background with BB
  • 238657 05/18/08 17:10 eschulte Gas Shed Air Flow Alarm
  • 238656 05/18/08 17:06 eschulte This should be labeled production
  • 238655 05/18/08 16:40 Kalyan Loop 3 temperature
  • 238654 05/18/08 16:33 eschulte Fixed
  • 238650 05/18/08 16:02 R. De Leo shift summary
  • 238643 05/18/08 14:56 khem set golden root file to 1293
  • 238638 05/18/08 14:11 Xiaohui FPP Hot wire in U1
  • 238636 05/18/08 14:05 ran elastic correlation BigBite
  • 238635 05/18/08 14:03 Reimer Target: 6 cm LH2
  • 238633 05/18/08 13:56 ran elastics on HRS
  • 238631 05/18/08 13:41 Guy Ron Ta Spectrum
  • 238630 05/18/08 13:36 Reimer Target: 6 cm Al Dummy
  • 238628 05/18/08 13:26 Reimer Target: Close JT 1 step
  • 238626 05/18/08 13:21 Reimer Target: Tantalum
  • 238624 05/18/08 13:17 brads Lost beam energy IOC
  • 238622 05/18/08 13:09 Reimer Target: Carbon
  • 238621 05/18/08 13:07 Guy Ron Dummy target vs. LH2 TargetThwe
  • 238616 05/18/08 12:54 reimer Target: Carbon w/hole
  • 238614 05/18/08 12:41 Guy Ron Asymmetry Histogram from FPP
  • 238613 05/18/08 12:04 adaq Radiation and Ozone snapshot
  • 238612 05/18/08 12:04 adaq Hall A Tools snapshot
  • 238611 05/18/08 12:04 adaq BigBox Magnet snapshot
  • 238610 05/18/08 12:03 adaq Beam Position and Current snapshot
  • 238609 05/18/08 12:03 adaq Beam Energy in A and C snapshot
  • 238607 05/18/08 11:57 Reimer Target: 6 cm Al Dummy
  • 238605 05/18/08 11:29 Guy Ron FPP Multiplicity
  • 238596 05/18/08 08:01 Solvignon Shift summary
  • 238586 05/18/08 04:57 S.Frullani hot wire FPP
  • 238585 05/18/08 04:52 brads FPP multiplicity reduction due to recentering beam?
  • 238582 05/18/08 04:43 brads Side effect of timing fix -- BB cosmics trigger (T7) will not work
  • 238581 05/18/08 04:11 Solvignon Spot++
  • 238579 05/18/08 04:03 adaq Radiation and Ozone snapshot
  • 238578 05/18/08 04:03 adaq Hall A Tools snapshot
  • 238577 05/18/08 04:03 adaq BigBox Magnet snapshot
  • 238576 05/18/08 04:03 adaq Beam Position and Current snapshot
  • 238575 05/18/08 04:03 adaq Beam Energy in A and C snapshot
  • 238573 05/18/08 04:02 cisbani Target status
  • 238572 05/18/08 04:01 cisbani Target stripcharts
  • 238571 05/18/08 03:51 xiaohui Fpp check
  • 238570 05/18/08 03:47 brads Things that should be investigated on Sunday
  • 238567 05/18/08 03:12 Xiaohui BB shower block hits
  • 238566 05/18/08 03:09 Kalyan BigBite preshower vs shower
  • 238565 05/18/08 03:07 xiaohui BB shower block 19.3 fixed
  • 238554 05/18/08 02:19 brads Solution(?) to the ADC issues and other Controlled Access activity
  • 238552 05/18/08 02:06 Solvignon Raster size is now 1.0x1.5
  • 238551 05/18/08 02:02 cisbani Target LH2 6cm (Loop3)
  • 238550 05/18/08 02:01 Solvignon Beam position update
  • 238539 05/18/08 01:18 cisbani Target Carbon w/hole
  • 238538 05/18/08 01:15 Solvignon Hall in beam permit
  • Saturday

  • 238530 05/17/08 23:57 Ramesh Shift summary
  • 238523 05/17/08 23:34 A. Camsonne AC filter for ADC
  • 238521 05/17/08 23:06 Nikos coincidence time
  • 238519 05/17/08 22:54 Ramesh Hall Access requested
  • 238510 05/17/08 21:07 Ramesh RF problem, beam stops
  • 238500 05/17/08 20:26 xiaohui Threshold change in BB shower and PS
  • 238498 05/17/08 20:18 Ramesh Moved to production Kin
  • 238493 05/17/08 20:11 Reimer Target: LH2 6cm
  • 238491 05/17/08 20:03 adaq Radiation and Ozone snapshot
  • 238490 05/17/08 20:03 adaq Hall A Tools snapshot
  • 238489 05/17/08 20:03 adaq BigBox Magnet snapshot
  • 238488 05/17/08 20:03 adaq Beam Position and Current snapshot
  • 238487 05/17/08 20:03 adaq Beam Energy in A and C snapshot
  • 238486 05/17/08 19:59 adaq Target Screens snapshot
  • 238482 05/17/08 19:33 Reimer Target: Empty
  • 238480 05/17/08 19:23 Ramesh Another test kinematics setting
  • 238474 05/17/08 18:56 Kalyan bigbite threshold changed
  • 238471 05/17/08 18:33 Ramesh Moved to new test kinematics setting
  • 238470 05/17/08 18:15 Guy Ron BB Problems
  • 238468 05/17/08 18:06 Kalyan Bigbite magnet current
  • 238464 05/17/08 17:32 Kalyan Bigbite magnet current changed
  • 238462 05/17/08 17:30 Reimer Target heat load OK
  • 238460 05/17/08 17:29 Ramesh Calibration/optimization cont.
  • 238459 05/17/08 17:21 Reimer Target PID and Beam compensation
  • 238455 05/17/08 17:10 brads BigBite magent current changes since 16:30
  • 238444 05/17/08 16:14 xiaohui HV changed for BB shower block 99(adc)
  • 238438 05/17/08 16:01 DeLeo run summary
  • 238429 05/17/08 14:41 Xiaohui Ped sub file added to coda
  • 238426 05/17/08 14:28 adaq Target Screens snapshot
  • 238413 05/17/08 11:11 ran dead time
  • 238403 05/17/08 09:30 eschulte 07:50 Beam Off Problem
  • 238401 05/17/08 08:18 eschulte Run Coordinator Summary 15 May Swing through 17 May Owl
  • 238400 05/17/08 08:09 Adaq Owl Shift summary
  • 238399 05/17/08 08:07 Norum Target status
  • 238397 05/17/08 07:55 eschulte This fixed the reconstruction problems
  • 238389 05/17/08 04:37 cisbani BPM stripchart
  • 238388 05/17/08 04:35 cisbani LHRS Magnet Chart
  • 238387 05/17/08 04:34 cisbani Hall A General Tools
  • 238384 05/17/08 04:03 adaq Radiation and Ozone snapshot
  • 238383 05/17/08 04:03 adaq Hall A Tools snapshot
  • 238382 05/17/08 04:02 adaq BigBox Magnet snapshot
  • 238381 05/17/08 04:02 adaq Beam Position and Current snapshot
  • 238380 05/17/08 04:02 adaq Beam Energy in A and C snapshot
  • 238379 05/17/08 04:02 S. Frullani replay software
  • 238378 05/17/08 04:02 adaq Target Screens snapshot
  • 238375 05/17/08 02:54 Adaq replay-software problem
  • 238372 05/17/08 01:57 cisbani memo: current BB HV setting file
  • 238371 05/17/08 01:47 Kalyan coincidence time plot
  • 238368 05/17/08 01:41 xiaohui FPP efficiency
  • 238367 05/17/08 01:33 Xiaohui LHRS Acceptance
  • 238366 05/17/08 00:55 ran coincidence time
  • 238365 05/17/08 00:53 gron FPP door position image
  • 238354 05/17/08 00:04 Solvignon Shift summary
  • 238353 05/17/08 00:03 eschulte Carbon Doors
  • Friday

  • 238348 05/16/08 23:41 adaq Radiation and Ozone snapshot
  • 238347 05/16/08 23:41 adaq Hall A Tools snapshot
  • 238346 05/16/08 23:41 adaq BigBox Magnet snapshot
  • 238343 05/16/08 23:37 adaq Beam Position and Current snapshot
  • 238342 05/16/08 23:37 adaq Beam Energy in A and C snapshot
  • 238340 05/16/08 23:29 Norum Target Aluminum
  • 238339 05/16/08 23:24 xiaohui BigBite Shower Block Hits
  • 238337 05/16/08 23:20 Xiaohui BB shower block map1
  • 238336 05/16/08 23:18 adaq Target Screens snapshot
  • 238331 05/16/08 22:51 Solvignon Spot++
  • 238330 05/16/08 22:38 Solvignon Beam positions: now good
  • 238318 05/16/08 22:01 xiaohui HRS acceptance
  • 238315 05/16/08 21:57 xiaohui FPP in good shape
  • 238314 05/16/08 21:55 Solvignon Left Q2 and Q3 were cycled
  • 238313 05/16/08 21:55 Xiaohui 6 cm Dummy Target Y reconstruction
  • 238312 05/16/08 21:54 Norum H2 target
  • 238311 05/16/08 21:53 Solvignon Bigbite magnet is now ON
  • 238309 05/16/08 21:51 Solvignon ROC8 is back
  • 238308 05/16/08 21:50 Norum Target status
  • 238306 05/16/08 21:40 Solvignon ROC8 problem
  • 238305 05/16/08 21:23 Solvignon Dipole NMR probe setup
  • 238304 05/16/08 21:22 Solvignon Hall A tools for production
  • 238303 05/16/08 21:16 Jack (Techs) Left Q3 polarity
  • 238302 05/16/08 21:05 Solvignon Q3 fixed. Back to beam permit
  • 238301 05/16/08 20:02 adaq Radiation and Ozone snapshot
  • 238300 05/16/08 20:02 adaq Hall A Tools snapshot
  • 238299 05/16/08 20:02 adaq Beam Position and Current snapshot
  • 238298 05/16/08 20:02 adaq Beam Energy in A and C snapshot
  • 238297 05/16/08 19:36 brads Followup() radiation monitors (beamline survey link)
  • 238296 05/16/08 19:34 Solvignon Update on left Q3 problem
  • 238295 05/16/08 18:08 Solvignon Moller done. Ask for controlled access
  • 238294 05/16/08 18:02 E.Chudakov Moller finished
  • 238292 05/16/08 17:47 brads Followup(^2) on radiation monitor issue
  • 238291 05/16/08 16:07 brads Followup on radiation monitor issue
  • 238290 05/16/08 16:03 jpchen Target screens
  • 238289 05/16/08 15:58 A. Camsonne BB Shower and PreShower thresholds added to epicRunstart
  • 238288 05/16/08 15:56 jpchen BCM strip charts
  • 238287 05/16/08 15:55 jpchen tools screen
  • 238286 05/16/08 15:52 brads Alarm handler bell off on hacsbc2
  • 238285 05/16/08 14:26 E.Chudakov Beam is back - ask for settings Moller again
  • 238284 05/16/08 14:20 Kalyan BigBite thresholds changed
  • 238283 05/16/08 13:11 E.Chudakov MCC reported a power glitch
  • 238282 05/16/08 13:05 E.Chudakov Starting a Moller measurement
  • 238281 05/16/08 13:02 brads Radiation monitor problem?
  • 238280 05/16/08 12:55 jpchen Big Bite magnet power off
  • 238279 05/16/08 12:52 jpchen day shift summary (partial)
  • 238274 05/16/08 12:39 jpchen run 1182 was 6 cm dummy target
  • 238272 05/16/08 12:34 jpchen BigBite HVs
  • 238270 05/16/08 12:13 adaq Hall A Tools snapshot
  • 238269 05/16/08 12:13 adaq Beam Position and Current snapshot
  • 238268 05/16/08 12:13 adaq Beam Energy in A and C snapshot
  • 238266 05/16/08 12:12 eschulte Hole-drilled target with 2x2 raster
  • 238264 05/16/08 12:03 eschulte Raster on Hole-drilled target
  • 238263 05/16/08 12:01 eschulte Raster off Hole-drilled Carbon target
  • 238223 05/16/08 08:03 cusanno Owl Shift Summary
  • 238221 05/16/08 07:40 Reimer Target: Empty
  • 238220 05/16/08 07:38 Reimer Target: LH2 6cm
  • 238216 05/16/08 06:30 Xiaohui FPP Performance
  • 238214 05/16/08 06:22 Xiaohui BB Shower Block Group
  • 238213 05/16/08 06:11 Reimer Target: JT Valve
  • 238212 05/16/08 06:10 Reimer Target: 6 cm Al Dummy
  • 238211 05/16/08 06:09 Xiaohui BB trigger
  • 238209 05/16/08 06:06 Xiaohui FPP HV Scan
  • 238205 05/16/08 05:41 cusanno Beam trip
  • 238197 05/16/08 04:52 cusanno Hall A Tools
  • 238192 05/16/08 04:28 cusanno BCM
  • 238187 05/16/08 04:13 adaq Hall A Tools snapshot
  • 238186 05/16/08 04:13 adaq Beam Position and Current snapshot
  • 238185 05/16/08 04:13 adaq Beam Energy in A and C snapshot
  • 238181 05/16/08 03:41 cusanno FPP HV scan
  • 238178 05/16/08 03:28 Reimer Target: LH2 6cm
  • 238177 05/16/08 03:04 cusanno Mozilla working on adaql8
  • 238176 05/16/08 03:04 cusanno Owl Shift Summary
  • 238170 05/16/08 02:44 cusanno Mozilla Firefox frozen on Shift Leader consolle
  • 238162 05/16/08 02:21 Reimer Target: Tantalum
  • 238159 05/16/08 02:14 Reimer Target: Carbon w/hole
  • 238158 05/16/08 02:11 cusanno Big Bite at 700 A
  • 238157 05/16/08 01:51 Reimer Target: Moved to Empty
  • 238155 05/16/08 01:37 cusanno Beam trips
  • 238154 05/16/08 01:32 brads Current monitor coeffs updated
  • 238151 05/16/08 01:11 cusanno Harp IH03A not working
  • 238143 05/16/08 00:10 eschulte "Center" Position
  • 238142 05/16/08 00:06 Seonho Choi Shift summary (Swing)
  • 238140 05/16/08 00:03 Seonho Choi Carbon hole and raster
  • Thursday

  • 238133 05/15/08 23:38 Seonho Choi Beam on target
  • 238132 05/15/08 23:36 LeRose Target Screens
  • 238131 05/15/08 22:58 Seonho Choi Hall A BCM not calibrated
  • 238130 05/15/08 22:56 Seonho Choi HARP Scan
  • 238129 05/15/08 22:40 Seonho Choi CW 10uA in the hall
  • 238128 05/15/08 22:13 Seonho Choi Pulsed beam into the hall
  • 238127 05/15/08 20:20 LeRose Target Encoder Positions
  • 238126 05/15/08 20:03 Seonho Choi L-HRS set for the first production kinematics
  • 238125 05/15/08 18:14 LeRose Closed JT Valve
  • 238124 05/15/08 17:23 eschulte Hall A is Now in Beam Permit
  • 238123 05/15/08 16:24 eschulte Hall A is now ready for sweep
  • 238121 05/15/08 15:54 Fansler Left Q3
  • 238120 05/15/08 15:53 Dezern Check list delivered
  • 238118 05/15/08 15:51 Dezern Checklist
  • 238116 05/15/08 15:35 jpchen LH2 target cooldown completed
  • 238113 05/15/08 15:17 eschulte Updated Checkout Plan
  • 238112 05/15/08 15:08 R. Michaels online deadtime modifications
  • 238105 05/15/08 13:45 dezern MCC has valve control
  • 238094 05/15/08 11:59 jpchen Cryotarget cool-down started.
  • 238083 05/15/08 11:37 Stevens/tech dipoles polarity reversal
  • 238080 05/15/08 11:32 Stevens/tech Left dipole phase rotation fault
  • 238051 05/15/08 10:35 meekins Hall A tgt pump/purge
  • 238039 05/15/08 10:11 eschulte E08-007 Start Up Checkout Plan
  • 238014 05/15/08 08:21 folts Hall Status
  • Wednesday

  • 237980 05/14/08 17:28 eschulte Status of Hall A on Wednesday 14 May 2008
  • 237939 05/14/08 10:55 eschulte Projected Target Collant
  • Tuesday

  • 237937 05/13/08 21:03 folts crane
  • 237929 05/13/08 17:42 eschulte Tuesday, May 13 RC Note
  • 237905 05/13/08 14:18 R. Michaels L-arm RT check
  • 237904 05/13/08 08:44 Jack (Techs) Right Q1 work yesterday (05-12-08)
  • Monday

  • 237896 05/12/08 19:37 V. Sulkosky RF Signal Added to L-HRS DAQ
  • 237895 05/12/08 19:05 brads Slow controls network switch problems fixed for now
  • 237889 05/12/08 14:04 R. Michaels T4 redefined. trigsetup defaults
  • 237888 05/12/08 13:59 atrig Trigger-Download
  • 237885 05/12/08 12:02 Jack Gas system restarted - had been down for more than a day
  • 237883 05/12/08 07:51 folts open Hall late entry
  • Sunday

  • 237880 05/11/08 19:58 Moffit Notes on EDTM setup for E08-007
  • 237876 05/11/08 15:50 R. Michaels EDTM mods
  • Saturday

  • 237865 05/10/08 14:55 eip BB electron stack
  • 237864 05/10/08 13:43 Moffit Trigger stuff into ROC8 TDCs.
  • Friday

  • 237861 05/09/08 19:36 Kalyan CODA TWOARM_FB config
  • 237850 05/09/08 17:30 LeRose Left Arm Dipole NMR studies Probe A vs Probe D
  • 237849 05/09/08 14:44 LeRose Left Arm Dipole NMR studies done
  • 237848 05/09/08 11:41 LeRose Left Arm Dipole NMR studies
  • Thursday

  • 237847 05/08/08 18:03 brads All guards installed on Hadron stack MWDCs
  • 237843 05/08/08 14:59 R. Michaels DAQ switchover for new experiment
  • 237839 05/08/08 14:21 brads Middle balloon removed, guards up on middle two MWDC windows
  • 237836 05/08/08 10:15 Higinbotham BigBite 48 degrees
  • 237835 05/08/08 10:15 Higinbotham 4He Gas Stopped and Disconnected
  • 237834 05/08/08 10:14 Higinbotham Safety: BigBite Electron Package Cable Installation
  • 237833 05/08/08 09:06 A. Camsonne Reminder TDC for trigger
  • 237832 05/08/08 09:06 A. Camsonne Reminder TDC for trigger
  • 237831 05/08/08 08:48 Lindgren BigBite and MWDC position
  • 237829 05/08/08 07:16 khem turn off bigbite hv and magnet, set momentum to 0 for both arms
  • 237819 05/08/08 05:55 Adaq Hall A Tools
  • 237810 05/08/08 04:09 adaq Hall A Tools snapshot
  • 237809 05/08/08 04:09 adaq Beam Position and Current snapshot
  • 237808 05/08/08 04:09 adaq Beam Energy in A and C snapshot
  • 237803 05/08/08 03:24 Nelyubin target screen
  • 237802 05/08/08 03:23 Nelyubin target snapshot
  • Wednesday

  • 237778 05/07/08 23:55 Seonho Choi Shift Summary (Swing)
  • 237769 05/07/08 22:25 D. McNulty Noticed that BPM4Ay is twice what it should be
  • 237766 05/07/08 22:11 D. McNulty Moving from Al. Dummy to loop1 LH2 (6cm)
  • 237758 05/07/08 20:54 D. McNulty Move target from LH2 to Al. dummy (6cm)
  • 237753 05/07/08 20:09 adaq Hall A Tools snapshot
  • 237752 05/07/08 20:09 adaq Beam Position and Current snapshot
  • 237751 05/07/08 20:09 adaq Beam Energy in A and C snapshot
  • 237739 05/07/08 18:35 Lindgren coin spectrum
  • 237730 05/07/08 17:00 Moffit Set golden run to 4960.
  • 237727 05/07/08 16:55 Moffit Proton Momentum with current E threshold settings
  • 237723 05/07/08 16:12 Moffit Changed E threshold for 4960
  • 237719 05/07/08 16:09 Ole Hansen Day shift summary
  • 237717 05/07/08 15:42 Tharanga Screen short
  • 237716 05/07/08 15:41 Tharanga Screen short
  • 237715 05/07/08 15:41 Tharanga screen short
  • 237714 05/07/08 15:38 Tharanga Screen shorts
  • 237712 05/07/08 15:37 Tharanga Target changes
  • 237711 05/07/08 15:36 Moffit Changed E threshold for 4958
  • 237709 05/07/08 14:53 Moffit Changed E threshold for 4957
  • 237702 05/07/08 14:12 Lindgren coin spectrum
  • 237701 05/07/08 14:10 Ole Hansen Current 7uA, raster on
  • 237698 05/07/08 14:01 Ole Hansen Beam energy lock on
  • 237697 05/07/08 13:58 Ole Hansen Hall A tools screenshot
  • 237694 05/07/08 13:45 Ole Hansen Kinematics Q, part III - production
  • 237688 05/07/08 13:41 Ole Hansen Current 1uA
  • 237684 05/07/08 13:33 Ole Hansen Hall A Tools
  • 237683 05/07/08 13:28 Ole Hansen Kinematics Q, part II - elastic H xsect
  • 237680 05/07/08 13:20 x. jiang FPP HV. turned master HV on to 1875,
  • 237677 05/07/08 12:59 Ole Hansen HRS-L quads cycled
  • 237676 05/07/08 12:50 Ole Hansen Kinematics Q part I - beam energy check
  • 237670 05/07/08 12:36 Ole Hansen Raster off
  • 237661 05/07/08 12:11 Ole Hansen Raster on
  • 237659 05/07/08 12:09 adaq Hall A Tools snapshot
  • 237658 05/07/08 12:09 adaq Beam Position and Current snapshot
  • 237657 05/07/08 12:09 adaq Beam Energy in A and C snapshot
  • 237654 05/07/08 11:58 Ole Hansen BB HV alarm
  • 237653 05/07/08 11:52 Ole Hansen Hall A tools screenshot
  • 237652 05/07/08 11:47 Ole Hansen Kinematics P, part III, more elastic hydrogen
  • 237647 05/07/08 11:39 xqian BB CHAMBER HV SCAN
  • 237646 05/07/08 11:39 xqian BB CHAMBER HV SCAN
  • 237644 05/07/08 11:38 Ole Hansen Current 5uA
  • 237640 05/07/08 11:31 xqian BB HV 1500 V
  • 237635 05/07/08 11:25 xqian HV 1700 V and 390 A
  • 237625 05/07/08 11:12 xqian HV 1650 BB
  • 237620 05/07/08 11:06 xqian HV BB is 1600 V
  • 237615 05/07/08 10:58 xqian HV screen at 1550 of BB
  • 237610 05/07/08 10:50 Ole Hansen BB MWDC HVs
  • 237607 05/07/08 10:48 Ole Hansen BB current 700A
  • 237602 05/07/08 10:43 Ole Hansen Raster off
  • 237601 05/07/08 10:38 x. jiang BB HV screen, 1 uA, LH2 target, 1500 volts.
  • 237598 05/07/08 10:35 Ole Hansen Current 1uA
  • 237595 05/07/08 10:32 x. jiang 5 uA, LH2 target. BB HV screen
  • 237592 05/07/08 10:21 Ole Hansen Kinematics P part II - elastic hydrogen
  • 237589 05/07/08 10:10 x. jiang BB HV screen, Carbon target 4.8 uA
  • 237586 05/07/08 10:06 Ole Hansen Raster on
  • 237581 05/07/08 09:45 Ole Hansen Current 5uA
  • 237578 05/07/08 09:37 Ole Hansen Kinematics P, part I
  • 237577 05/07/08 09:36 Higinbotham BigBite Measurements (late entry)
  • 237576 05/07/08 09:31 Ole Hansen Current 3uA
  • 237571 05/07/08 09:23 Ole Hansen Raster off
  • 237564 05/07/08 09:05 Ole Hansen Current 5uA
  • 237559 05/07/08 08:59 Ole Hansen Raster on
  • 237556 05/07/08 08:56 Tharanga screen short
  • 237553 05/07/08 08:51 Ole Hansen C target, 7uA, raster off
  • 237552 05/07/08 08:43 Ole Hansen Change to Kinematics O (part IV) for C checks
  • 237547 05/07/08 08:02 cisbani owl shift summary
  • 237537 05/07/08 06:30 cisbani BB HV Trip
  • 237531 05/07/08 05:34 cisbani Target strips
  • 237530 05/07/08 05:34 cisbani Target Status
  • 237523 05/07/08 04:43 cisbani BB HV Slot 14 correct!
  • 237522 05/07/08 04:39 cisbani BB HV Slot 14
  • 237517 05/07/08 04:09 adaq Hall A Tools snapshot
  • 237516 05/07/08 04:09 adaq Beam Position and Current snapshot
  • 237515 05/07/08 04:09 adaq Beam Energy in A and C snapshot
  • 237512 05/07/08 03:49 cisbani Correct Left Arm Angle in run 4914
  • 237505 05/07/08 03:03 cisbani Left Arm Monitor Blank, IOC Rebooted
  • 237492 05/07/08 01:23 Adaq Beam position unstable during run 4905
  • 237480 05/07/08 00:04 Seonho Choi Shift summary (swing)
  • Tuesday

  • 237477 05/06/08 23:51 cusanno run 4901 and run 4902 are calibration runs
  • 237475 05/06/08 23:46 Armstrong Target: move back to 6 cm lH2 target
  • 237470 05/06/08 22:35 Armstrong Target: move to 6cm dummy
  • 237464 05/06/08 21:35 Seonho Choi BPM/BCM Screen Capture
  • 237463 05/06/08 21:34 Seonho Choi Beam energy graph capture
  • 237462 05/06/08 21:34 Seonho Choi Radiation/Ozone Monitor Capture
  • 237461 05/06/08 21:33 Seonho Choi Hall A General Tools Capture
  • 237450 05/06/08 20:09 adaq Hall A Tools snapshot
  • 237449 05/06/08 20:09 adaq Beam Position and Current snapshot
  • 237448 05/06/08 20:08 adaq Beam Energy in A and C snapshot
  • 237437 05/06/08 19:09 Armstrong Target: screen shots
  • 237422 05/06/08 17:05 Jack Proportional tubes added during access today
  • 237419 05/06/08 16:42 Armstrong go to 7 µA
  • 237416 05/06/08 16:02 x. jiang day shift summary
  • 237402 05/06/08 14:09 hyao target moved from carbon to Loop1: 6cm (H2)
  • 237393 05/06/08 13:33 Adaq TA target run(4871) analysis result
  • 237390 05/06/08 13:25 X. Jiang BigBite magnet screen
  • 237389 05/06/08 13:23 hyao target moved from tantalum to carbon
  • 237388 05/06/08 13:23 X. Jiang Hall A tools screen.
  • 237380 05/06/08 12:23 S. Marrone target change
  • 237377 05/06/08 12:08 adaq Hall A Tools snapshot
  • 237376 05/06/08 12:08 adaq Beam Position and Current snapshot
  • 237375 05/06/08 12:08 adaq Beam Energy in A and C snapshot
  • 237364 05/06/08 10:59 Ramesh New momentum setting; Q2, Q3 cycled
  • 237363 05/06/08 10:50 brads He bag topped off (should be good through to Thursday)
  • 237359 05/06/08 10:44 Higinbotham Cosmic Runs with different prescale
  • 237357 05/06/08 09:03 x. jiang Beam off. Accelerator beam study. Four MCC hours.
  • 237355 05/06/08 08:57 S. Marrone Spectra Deadtime and ToF run 4860
  • 237354 05/06/08 08:47 x. jiang spot++ beam position check
  • 237353 05/06/08 08:46 X. Jiang Beam energy strip chart
  • 237352 05/06/08 08:46 X. Jiang BPM strip chart.
  • 237351 05/06/08 08:45 X. Jiang Hall A tools and beam current screenshot
  • 237350 05/06/08 08:44 x. jiang BigBite HV screen shot, 5.0 uA beam.
  • 237343 05/06/08 08:06 A. Camsonne Owl shift summary
  • 237338 05/06/08 07:40 A. Camsone more plots for run 4801
  • 237336 05/06/08 07:27 A. Camsonne Run 4801 Carbon high deadtime
  • 237331 05/06/08 07:08 cisbani Target to H2 (Loop1 6cm)
  • 237325 05/06/08 06:26 A. Camsonne BigBite MWDC HV
  • 237323 05/06/08 06:23 cisbani Target in Dummy position
  • 237322 05/06/08 06:19 A. Camsonne Beam energy
  • 237321 05/06/08 06:16 A. Camsonne BCM/BPM
  • 237319 05/06/08 06:14 A. Camsonne Hall A tools
  • 237317 05/06/08 06:05 cisbani Target status
  • 237299 05/06/08 04:08 adaq Hall A Tools snapshot
  • 237298 05/06/08 04:08 adaq Beam Position and Current snapshot
  • 237297 05/06/08 04:08 adaq Beam Energy in A and C snapshot
  • Monday

  • 237262 05/05/08 23:59 Seonho Choi Shift summary (swing)
  • 237249 05/05/08 22:05 Seonho Choi Ozone/Rad Mon Strip Chart
  • 237248 05/05/08 22:05 Seonho Choi Hall-A Tools Screen Capture
  • 237247 05/05/08 22:04 Seonho Choi BPM/BCM Strip Chart
  • 237246 05/05/08 22:03 Seonho Choi Beam Energy Strip Chart
  • 237245 05/05/08 22:01 cusanno Target Screenshot
  • 237226 05/05/08 20:08 adaq Hall A Tools snapshot
  • 237225 05/05/08 20:08 adaq Beam Position and Current snapshot
  • 237224 05/05/08 20:08 adaq Beam Energy in A and C snapshot
  • 237217 05/05/08 19:27 Moffit Wien angle changed, just before run 4808
  • 237209 05/05/08 18:19 Seonho Choi Beam off to adjust orbit
  • 237201 05/05/08 16:49 Moffit Disabled plots in physics replay display
  • 237198 05/05/08 16:33 jinhuang Bigbite Lib update
  • 237194 05/05/08 16:05 LeRose Shift Summary (Day)
  • 237189 05/05/08 15:39 LeRose Beam Back
  • 237188 05/05/08 14:59 LeRose BigBite HV
  • 237187 05/05/08 14:58 LeRose BigBite Current
  • 237186 05/05/08 14:58 Armstrong Target: screen shots
  • 237185 05/05/08 14:54 LeRose BPM/BCM/Energy
  • 237184 05/05/08 14:52 LeRose BB HV Screen
  • 237182 05/05/08 14:31 Moffit Changed/recompiled BBNormAna and replay script
  • 237181 05/05/08 14:22 brads Controlled access activity (late entry)
  • 237180 05/05/08 14:17 LeRose Left Arm: 2.045 GeV/c
  • 237179 05/05/08 14:13 LeRose No Beam
  • 237175 05/05/08 13:54 Armstrong Target: moved to 6 cm lH2
  • 237174 05/05/08 13:53 LeRose Left Arm: 2.18 GeV/c
  • 237171 05/05/08 13:41 LeRose Deadtime ~50%
  • 237166 05/05/08 13:20 LeRose Beam Back
  • 237165 05/05/08 13:15 LeRose No Beam for a few minutes
  • 237164 05/05/08 13:12 Armstrong Target: move to Carbon
  • 237161 05/05/08 12:08 adaq Hall A Tools snapshot
  • 237160 05/05/08 12:08 adaq Beam Position and Current snapshot
  • 237159 05/05/08 12:08 adaq Beam Energy in A and C snapshot
  • 237158 05/05/08 11:46 Armstrong Target: to Ta
  • 237157 05/05/08 11:17 LeRose BB HV restored
  • 237156 05/05/08 11:11 LeRose Harp Scan / Controlled Access
  • 237155 05/05/08 11:04 jpchen LD2 loop temperature set point changed to 22K
  • 237154 05/05/08 11:02 LeRose BB HV trip
  • 237153 05/05/08 10:52 Armstrong Target: move to empty
  • 237152 05/05/08 10:49 Armstrong Target: alarms unsilenced
  • 237151 05/05/08 10:47 jpchen Adjustment to loop 1/loop 2.
  • 237150 05/05/08 10:46 LeRose CW beam
  • 237149 05/05/08 09:54 LeRose Left HRS to 2.323 GeV/c
  • 237148 05/05/08 09:49 LeRose Pulsed Beam
  • 237147 05/05/08 09:14 LeRose Pass Change
  • 237146 05/05/08 08:49 Armstrong Target: loop 1 fan back to normal
  • 237139 05/05/08 08:10 Armstrong Target: screen shots
  • 237137 05/05/08 07:56 R. Michaels owl shift summary
  • 237133 05/05/08 07:43 R. Michaels alum. dummy runs
  • 237123 05/05/08 06:23 R. Michaels target alarm
  • 237110 05/05/08 04:41 cusanno BigBite HV screenshot
  • 237109 05/05/08 04:41 cusanno Beam Energy screenshot
  • 237108 05/05/08 04:40 cusanno BPM/BCM stripchart
  • 237107 05/05/08 04:38 cusanno Hall A Tools
  • 237104 05/05/08 04:25 R. Michaels tools screen
  • 237103 05/05/08 04:25 R. Michaels BB HV screen
  • 237102 05/05/08 04:23 R. Michaels BPM stripchart
  • 237101 05/05/08 04:23 R. Michaels energy stripchart
  • 237098 05/05/08 04:08 adaq Hall A Tools snapshot
  • 237097 05/05/08 04:08 adaq Beam Position and Current snapshot
  • 237096 05/05/08 04:08 adaq Beam Energy in A and C snapshot
  • 237091 05/05/08 03:24 R. Michaels beam current
  • 237080 05/05/08 01:59 cusanno Beam current 5.8 uA
  • 237078 05/05/08 01:57 R. Michaels adjust I_beam to 6 uA.
  • 237073 05/05/08 01:16 R. Michaels reduce current to 5.5 uA
  • 237069 05/05/08 00:57 jpchen Loop 1 (LH2) fan not well
  • 237066 05/05/08 00:39 R. Michaels beam returns
  • 237065 05/05/08 00:38 R. Michaels Target Fan speed a bit low
  • 237064 05/05/08 00:24 R. Michaels No beam (MCC calls)
  • 237062 05/05/08 00:18 R. Michaels beam back, BB HV trip/reset
  • 237060 05/05/08 00:07 R. Michaels status - kine H (magnets)
  • 237058 05/05/08 00:05 X. Jiang swing shift summary
  • 237057 05/05/08 00:04 Kalyan back to beam permit
  • Sunday

  • 237056 05/04/08 23:40 Kalyan going to controlled access
  • 237054 05/04/08 23:18 x. jiang BB HV trip: bad power supply 24 v, etc.
  • 237046 05/04/08 22:31 Kalyan Target screenshots
  • 237045 05/04/08 22:26 Moffit Golden set to 4761
  • 237034 05/04/08 21:25 Kalyan Deadtime around 20%
  • 237033 05/04/08 21:14 X. Jiang BB chamber HV, at 6.8uA beam
  • 237032 05/04/08 21:13 x. jiang BigBite magnet control screen
  • 237030 05/04/08 21:10 x. jiang BPM strip chart
  • 237029 05/04/08 21:07 x. jiang BB HV tripped, all channels.
  • 237027 05/04/08 21:00 x. jiang run 4761, 6.7 uA beam, raster on
  • 237023 05/04/08 20:18 Moffit Golden set to 4758 for kin H
  • 237021 05/04/08 20:08 adaq Hall A Tools snapshot
  • 237020 05/04/08 20:08 adaq Beam Position and Current snapshot
  • 237019 05/04/08 20:08 adaq Beam Energy in A and C snapshot
  • 237016 05/04/08 19:26 Kalyan Target is loop 1 hydrogen
  • 237015 05/04/08 19:19 x. jiang left HRS momentum set to 0.960
  • 237012 05/04/08 18:50 X. Jiang BigBite HV tripped, reset.
  • 237007 05/04/08 18:29 x. jiang left-HRS momentum to 1105 MeV/c
  • 237001 05/04/08 17:31 Kalyan Target moved to Ta
  • 236994 05/04/08 16:47 x. jiang Hall A tool screen, left arm angle changed to 20.5
  • 236990 05/04/08 16:38 Lindgren Shift summary
  • 236950 05/04/08 14:57 cusanno Carbon target in
  • 236945 05/04/08 14:29 cusanno Target Alarm Handler fixed
  • 236942 05/04/08 14:21 Lindgren Coin Spectrum
  • 236933 05/04/08 13:13 Lindgren Coin spectrum
  • 236932 05/04/08 13:08 Lindgren MWDC HV at 2 ua Current
  • 236929 05/04/08 12:36 cusanno Target screenshots
  • 236928 05/04/08 12:32 Lindgren Beam Current strip chart
  • 236925 05/04/08 12:24 cusanno Beam current in the start-of-run entries
  • 236924 05/04/08 12:08 adaq Hall A Tools snapshot
  • 236923 05/04/08 12:08 adaq Beam Position and Current snapshot
  • 236922 05/04/08 12:08 adaq Beam Energy in A and C snapshot
  • 236920 05/04/08 12:02 Lindgren beam current decrease
  • 236916 05/04/08 11:42 Lindgren Correction for Entry 236886,236894,236903,
  • 236913 05/04/08 11:27 Lindgren Hall A Tools and Current stripchart
  • 236912 05/04/08 11:25 Lindgren bigbite scalers
  • 236911 05/04/08 11:21 Lindgren MWDC HV
  • 236910 05/04/08 11:19 Lindgren Lower beam current test
  • 236903 05/04/08 10:33 Lindgren S3 HV Crashed
  • 236896 05/04/08 09:59 cusanno Target Alarm Handler
  • 236895 05/04/08 09:58 cusanno Target Pressure and Temperature
  • 236894 05/04/08 09:54 Lindgren Dead Time
  • 236888 05/04/08 09:34 Lindgren Increase current
  • 236886 05/04/08 09:32 Lindgren Target Change
  • 236879 05/04/08 08:39 cusanno LH2 Loop 1 (w/o coll) in
  • 236875 05/04/08 08:17 M.Shabestari Owl Shift Summary
  • 236871 05/04/08 08:01 M.Shabestari Target screens
  • 236868 05/04/08 07:50 M.Shabestari BigBite HV snapshot
  • 236863 05/04/08 07:24 M.Shabestari Hall A Tools snapshot
  • 236860 05/04/08 07:09 M.Shabestari Radiation stripchart
  • 236857 05/04/08 07:00 M.Shabestari BCM and BPM snapshots
  • 236854 05/04/08 06:44 M.Shabestari Target moved to LH2 with collimator
  • 236853 05/04/08 05:10 M.Shabestari Target moved to empty
  • 236852 05/04/08 04:24 M.Shabestari No beam due to RF problem
  • 236851 05/04/08 04:08 adaq Hall A Tools snapshot
  • 236850 05/04/08 04:08 adaq Beam Position and Current snapshot
  • 236849 05/04/08 04:08 adaq Beam Energy in A and C snapshot
  • 236836 05/04/08 01:49 M.Shabestari Beam energy stripchart
  • 236833 05/04/08 01:37 M.Shabestari HV on S1 and S14 tripped
  • 236829 05/04/08 01:24 M.Shabestari Target moved to LH2 with collimator
  • 236825 05/04/08 00:48 M.Shabestari BigBite HV snapshot
  • 236824 05/04/08 00:45 M.Shabestari BCM and BPM stripcharts
  • 236823 05/04/08 00:44 M.Shabestari Hall A Tools snapshot
  • 236818 05/04/08 00:26 M.Shabestari Target moved to Al
  • 236812 05/04/08 00:01 cusanno Swing Shift Summary
  • Saturday

  • 236793 05/03/08 22:02 Armstrong Target: loop 2 is just fine.
  • 236789 05/03/08 21:39 cusanno Ozone stripchart
  • 236787 05/03/08 21:36 cusanno Beam off for a while
  • 236786 05/03/08 21:31 Armstrong Target: figure for loop2 oscillations
  • 236785 05/03/08 21:27 Armstrong Target: PID on loop 2 oscillates
  • 236782 05/03/08 21:13 Moffit Left Dipole NMR not locked for run 4675
  • 236781 05/03/08 21:08 cusanno BigBite Wire Chamber trip
  • 236778 05/03/08 20:58 cusanno Big Bite HV slot 14
  • 236777 05/03/08 20:56 cusanno Beam Energy stripchart
  • 236776 05/03/08 20:55 cusanno BPM/BCM stripchart
  • 236775 05/03/08 20:53 cusanno Hall A Tools
  • 236774 05/03/08 20:51 cusanno Production restored
  • 236772 05/03/08 20:35 cusanno Big Bite MWPC tripped again
  • 236769 05/03/08 20:25 cusanno Big-Bite HV restarted
  • 236768 05/03/08 20:08 adaq Hall A Tools snapshot
  • 236767 05/03/08 20:08 adaq Beam Position and Current snapshot
  • 236766 05/03/08 20:08 adaq Beam Energy in A and C snapshot
  • 236765 05/03/08 20:07 cusanno Trip of Big-Bite MP Wire Chamber
  • 236763 05/03/08 20:04 cusanno Big Bite Angle should be 48 deg
  • 236761 05/03/08 19:47 cusanno HRS-Left Momentum 997 MeV
  • 236758 05/03/08 19:29 cusanno Dead Time is back < 20% on Elastic H2
  • 236756 05/03/08 19:19 Armstrong Target to 6 cm hydrogen
  • 236755 05/03/08 19:14 cusanno HRS Left Momentum 1.147 GeV
  • 236752 05/03/08 18:49 Armstrong Move to Carbon target; deadtime
  • 236751 05/03/08 18:47 cusanno Carbon target in
  • 236749 05/03/08 18:39 Armstrong Target status
  • 236747 05/03/08 18:35 cusanno Beam back
  • 236745 05/03/08 18:26 cusanno Beam trip
  • 236744 05/03/08 18:25 cusanno Dead Time
  • 236742 05/03/08 17:58 cusanno Ta target in
  • 236741 05/03/08 17:56 Armstrong Target: fans restarted
  • 236740 05/03/08 17:50 cusanno Beam permit
  • 236739 05/03/08 17:49 cusanno target fan back
  • 236738 05/03/08 17:34 cusanno JP went in Controlled Access
  • 236737 05/03/08 17:28 cusanno BB at 31.1 deg
  • 236736 05/03/08 17:28 cusanno BB at 31.1 deg
  • 236735 05/03/08 17:28 cusanno BB at 31.1 deg
  • 236734 05/03/08 17:00 Ramesh Preparation for Kin F
  • 236733 05/03/08 16:55 cusanno HRS moved to 14.5 deg
  • 236732 05/03/08 16:44 Armstrong controlled access
  • 236728 05/03/08 16:16 Armstrong move to Dummy target
  • 236724 05/03/08 16:08 Lindgren Day Shift Summary
  • 236722 05/03/08 16:01 Armstrong Target status
  • 236717 05/03/08 15:23 Moffit Re: Strangeness
  • 236712 05/03/08 14:41 xqian Elastic proton strip -- Optics
  • 236708 05/03/08 14:32 Moffit Golden run 4648 re-replayed.
  • 236706 05/03/08 14:06 Moffit updated db_run.dat for runs 4621-4660
  • 236705 05/03/08 13:59 Moffit updated mysql database for runs 4621-4660
  • 236702 05/03/08 13:25 Lindgren Hall A Toola and Current strip
  • 236701 05/03/08 13:23 Lindgren BPM/BCM
  • 236700 05/03/08 13:21 Lindgren MWDC HV
  • 236699 05/03/08 13:19 Lindgren Beam down
  • 236696 05/03/08 12:55 Lindgren Strangeness
  • 236693 05/03/08 12:32 ycwang ycwang
  • 236690 05/03/08 12:13 ycwang Target temperature goes down, and keeps 18.15, the fan does not work.
  • 236689 05/03/08 12:12 ycwang Target temperature becomes lower, the fan does not work.
  • 236688 05/03/08 12:08 adaq Hall A Tools snapshot
  • 236687 05/03/08 12:08 adaq Beam Position and Current snapshot
  • 236686 05/03/08 12:08 adaq Beam Energy in A and C snapshot
  • 236683 05/03/08 11:51 ycwang Target temperature becomes lower
  • 236682 05/03/08 11:49 ycwang Target temperature becomes lower
  • 236679 05/03/08 11:12 Lindgren Deadtime studies
  • 236676 05/03/08 11:04 A. camsonne Long replay for run 4627 and 4626 deuterium
  • 236673 05/03/08 10:55 Lindgren Current increase/Deadtime
  • 236672 05/03/08 10:47 A. camsonne Long replay for run 4627 and 4626 deuterium
  • 236667 05/03/08 10:21 A. Camsonne Corrected kinematics for run 4644 4646 4647
  • 236658 05/03/08 09:05 A. Camsonne updated kinematic name to B in currentKineName
  • 236656 05/03/08 08:56 Lindgren Going to Production Runing
  • 236654 05/03/08 08:44 A. Camsonne Owl shift summary
  • 236653 05/03/08 08:44 A. Camsonne High deadtime for run 4645 C
  • 236652 05/03/08 08:41 Lindgren Run 4647 Starting
  • 236651 05/03/08 08:40 Lindgren HV Crashed
  • 236647 05/03/08 08:30 R. Lindgren Starting day Shift
  • 236646 05/03/08 08:26 ycwang after moving target to H2
  • 236643 05/03/08 08:09 ycwang target screens
  • 236642 05/03/08 08:07 A. Camsonne Left Arm momentum set to 1.1338 GeV/c
  • 236639 05/03/08 07:55 A. Camsonne Beam energy chart
  • 236638 05/03/08 07:52 A. Camsonne Bigbite HV Slot 14
  • 236637 05/03/08 07:50 A. Camsonne Ozone radiation charts
  • 236636 05/03/08 07:49 A. Camsonne BCM/BPM chart
  • 236633 05/03/08 07:48 A. Camsonne BCM chart
  • 236632 05/03/08 07:45 A. Camsonne Moving to C target
  • 236629 05/03/08 07:29 A. Camsonne Left arm angle set to 16.5 degrees
  • 236626 05/03/08 07:20 A. Camsonne Target screen
  • 236625 05/03/08 07:20 A. Camsonne Target screens with no beam
  • 236624 05/03/08 07:16 A. Camsonne Target moved to Ta target
  • 236623 05/03/08 07:12 A. Camsonne Magnets setting after Q2 and Q3 cycling for Kin B
  • 236620 05/03/08 06:52 A. Camsonne Hall A tools
  • 236617 05/03/08 06:45 A.Camsonne Loop 2 screens and JT valve
  • 236616 05/03/08 06:43 A. Camsonne Loop 1 JT valve and screens
  • 236615 05/03/08 06:42 A. Camsonne Running condition for loop 1
  • 236612 05/03/08 06:38 A. Camsonne Left HRS Sieve out
  • 236609 05/03/08 06:32 A. Camsonne Oscillations on Loop 2
  • 236606 05/03/08 06:09 A. Camsonne BigBite Current to 390 A
  • 236601 05/03/08 05:57 A. Camsonne BigBite High Voltage trip
  • 236595 05/03/08 05:24 A. Camsonne Bigbite current set to 600 A
  • 236591 05/03/08 04:41 A. Camsonne BigBite Current set to 500 A
  • 236588 05/03/08 04:08 adaq Hall A Tools snapshot
  • 236587 05/03/08 04:08 adaq Beam Position and Current snapshot
  • 236586 05/03/08 04:07 adaq Beam Energy in A and C snapshot
  • 236585 05/03/08 03:17 A. Camsonne increased current to 5uA
  • 236582 05/03/08 03:03 A. Camsonne Left arm Sieve in
  • 236581 05/03/08 03:01 A. Camsonne BigBite surrent set to 200 A
  • 236578 05/03/08 02:42 A. Camsonne Loop 1 fan speed set to 12 %
  • 236577 05/03/08 02:38 A. Camsonne Reduced Loop 2 by one step increased Loop 1 JT by one step
  • 236576 05/03/08 02:36 A. Camsonne Fan speed Loop 1 to 11%
  • 236573 05/03/08 02:31 A. Camsonne Reduced Loop 2 JT valve one more step
  • 236572 05/03/08 02:29 A. Camsonne Loop 1 fan speed set to 10 % and opened JT valve by 0.3
  • 236571 05/03/08 02:23 A. Camsonne Run 4635 LH2 with various fan speed and coolant flow
  • 236570 05/03/08 02:20 A. Camsonne Reduced Loop2 JT valve one more step
  • 236569 05/03/08 02:19 A. Camsonne Increased loop 1 fan speed to 9%
  • 236568 05/03/08 02:11 A. Camsonne Target chart Loop 1
  • 236567 05/03/08 02:10 A. Camsonne Closed Loop 2 JT valve by 0.3
  • 236564 05/03/08 02:03 A. Camsonne Moving to Loop 1 and starting reducing L2 fan speed
  • 236559 05/03/08 01:36 A. Camsonne Target screens LD2
  • 236558 05/03/08 01:35 A. Camsonne Target screens LD2
  • 236555 05/03/08 01:28 A. Camsonne Moved back to LD2 target
  • 236554 05/03/08 01:10 A. Camsonne moving to Empty target for beam tune
  • 236553 05/03/08 00:07 Moffit A few changes to replay.
  • 236552 05/03/08 00:06 Yi Zhang Swing shift summary
  • Friday

  • 236538 05/02/08 21:29 Y. Zhang BB sieve slit
  • 236535 05/02/08 20:53 Tharanga Target
  • 236534 05/02/08 20:52 Tharanga target
  • 236529 05/02/08 20:07 adaq Hall A Tools snapshot
  • 236528 05/02/08 20:07 adaq Beam Position and Current snapshot
  • 236527 05/02/08 20:07 adaq Beam Energy in A and C snapshot
  • 236524 05/02/08 19:59 Moffit and C Added BBNormAna to replay.
  • 236519 05/02/08 18:52 A.Rakhman LD2 SCREEN SHOT
  • 236506 05/02/08 17:56 A.Rakhman Target screen shot, calibration, carbon target
  • 236503 05/02/08 17:24 jpchen LD2 target ready for beam
  • 236497 05/02/08 16:02 bogdanw shift summary
  • 236483 05/02/08 14:08 bogdanw HRS-L for P33
  • 236473 05/02/08 12:07 adaq Hall A Tools snapshot
  • 236472 05/02/08 12:07 adaq Beam Position and Current snapshot
  • 236471 05/02/08 12:07 adaq Beam Energy in A and C snapshot
  • 236469 05/02/08 11:52 bogdanw set for kin M
  • 236446 05/02/08 09:25 R. Michaels mini-sign-check
  • 236438 05/02/08 08:32 bogdanw beam energy
  • 236437 05/02/08 08:32 bogdanw BPM and ozone
  • 236436 05/02/08 08:30 bogdanw tools snapshot
  • 236428 05/02/08 07:59 Nilanga Shift Summary
  • 236419 05/02/08 07:05 Vladimir Nel snapshot of target screen
  • 236408 05/02/08 05:39 Nilanga BPM strip
  • 236391 05/02/08 04:07 adaq Hall A Tools snapshot
  • 236390 05/02/08 04:07 adaq Beam Position and Current snapshot
  • 236389 05/02/08 04:07 adaq Beam Energy in A and C snapshot
  • 236370 05/02/08 02:12 Nilanga MWDC HV
  • 236369 05/02/08 02:11 Vladimir Nel snapshot target screen
  • 236368 05/02/08 02:10 Nilanga BCM Strip
  • 236367 05/02/08 02:09 Vladimir Nel snapshot target screen
  • 236366 05/02/08 02:08 Nilanga Hall A tools
  • 236345 05/02/08 00:12 LeRose Shift Summary (swing) cont'd
  • 236342 05/02/08 00:07 LeRose Shift Summary (swing)
  • Thursday

  • 236338 05/01/08 23:53 jinhuang replay script update: Normalization Analysis Module Added
  • 236337 05/01/08 23:53 jinhuang replay script update: Normalization Analysis Module Added
  • 236329 05/01/08 23:10 LeRose BB HV Trip
  • 236323 05/01/08 22:49 Chiran TARGET STATUS
  • 236320 05/01/08 22:41 brads Real deadtime for run 4540 was 13%
  • 236299 05/01/08 20:52 Chiranjib TARGET ------>>. L1
  • 236294 05/01/08 20:09 Yi Qiang Negative Dead Time
  • 236293 05/01/08 20:07 adaq Hall A Tools snapshot
  • 236292 05/01/08 20:07 adaq Beam Position and Current snapshot
  • 236291 05/01/08 20:07 adaq Beam Energy in A and C snapshot
  • 236288 05/01/08 20:03 Chiranjib TARGET---->> Al
  • 236248 05/01/08 16:32 LeRose BB HV Trip
  • 236246 05/01/08 16:24 Chiranjib TARGET STATUS
  • 236240 05/01/08 16:00 gilman day shift summary
  • 236239 05/01/08 15:59 gilman lowering HRS left momentum
  • 236236 05/01/08 15:50 xqian tracking efficiency and hitting efficiency
  • 236232 05/01/08 15:45 gilman beam trip related BB HV trip
  • 236231 05/01/08 15:36 gilman target / momentum change
  • 236226 05/01/08 15:14 Norum Target status
  • 236220 05/01/08 14:54 gilman setting hrs left for tantalum
  • 236219 05/01/08 14:36 gilman power back up
  • 236218 05/01/08 14:02 gilman BB HV
  • 236217 05/01/08 14:01 gilman escorted access
  • 236210 05/01/08 13:08 Ramesh Kin C production finished
  • 236200 05/01/08 12:07 adaq Hall A Tools snapshot
  • 236199 05/01/08 12:07 adaq Beam Position and Current snapshot
  • 236198 05/01/08 12:07 adaq Beam Energy in A and C snapshot
  • 236182 05/01/08 10:14 gilman beam off about 10 minutes
  • 236179 05/01/08 09:59 gilman BB hv trip
  • 236162 05/01/08 08:09 M.Shabestari Owl Shift Summary
  • 236159 05/01/08 07:57 M.Shabestari BigBite trigger planes single rates
  • 236156 05/01/08 07:53 M.Shabestari BigBite HV snapshot
  • 236154 05/01/08 07:50 M.Shabestari Hall A Tools snapshot
  • 236144 05/01/08 06:55 M.Shabestari BCM, BPM, and Beam energy stripcharts
  • 236137 05/01/08 06:15 M. Shabestar HV on S2 and S14 tripped
  • 236128 05/01/08 05:45 M.Shabestari HV tripped
  • 236121 05/01/08 05:22 M.Shabestari Radiation stripchart
  • 236118 05/01/08 05:09 M.Shabestari HV tripped due to beam trip
  • 236111 05/01/08 04:51 M.Shabestari Target screens
  • 236106 05/01/08 04:19 M.Shabestari HV tripped again
  • 236101 05/01/08 04:07 adaq Hall A Tools snapshot
  • 236100 05/01/08 04:07 adaq Beam Position and Current snapshot
  • 236099 05/01/08 04:07 adaq Beam Energy in A and C snapshot
  • 236088 05/01/08 03:00 M.Shabestari BigBite HV snapshot
  • 236085 05/01/08 02:54 Aniol tgt charts
  • 236084 05/01/08 02:53 M.Shabestari Beam energy stripchart
  • 236081 05/01/08 02:39 M.Shabestari HV tripped
  • 236076 05/01/08 02:17 M.Shabestari HV tripped
  • 236065 05/01/08 01:34 M.Shabestari Rebooted ROCs
  • 236057 05/01/08 00:36 M.Shabestari HV on S4, S5, and S14 tripped
  • 236054 05/01/08 00:31 M.Shabestari BigBite HV snapshot
  • 236051 05/01/08 00:27 M.Shabestari BCM and BPM stripcharts
  • 236050 05/01/08 00:23 M.Shabestari Hall A Tools snapshot
  • 236045 05/01/08 00:10 Norum Shift summary Part 2